### 18 Mar 2026 v2

# General Bounds on Functionals of the Lifetime under Life Table Constraints in a Joint Actuarial-Financial Framework

Abstract
In life insurance, life tables are used to estimate the survival distribution of individuals from a given population. However, these tables only provide survival probabilities at integer ages but no information about the distribution of deaths between two consecutive integer values. This incompleteness is particularly relevant for modern insurance products such as variable annuities, whose payoffs depend jointly on lifetime uncertainty and financial market performance. The valuation of such contracts must therefore be carried out in a joint actuarial-financial framework, as their values depend not only on the full information about mortality rates but also on the interaction between mortality risk, asset dynamics, and embedded guarantees. One frequent solution to this incompleteness is to postulate fractional age assumptions or mortality rate models, but it turns out that the results of the computations strongly depend on these assumptions, which makes it difficult to generalize them. We hence derive upper and lower bounds of hybrid functionals of the lifetime with respect to mortality rates, which are compatible with the observed life table at integer ages and the given financial market. We derive two sets of results under distinct assumptions. In the first, we assume that each mortality trajectory is almost surely consistent with all the given one-year survival probabilities from the table. In the second, we consider a relaxed formulation that allows for deviations of the mortality rates while still being consistent in expectation with the given one-year reference survival probabilities. These distinct yet complementary approaches provide a new robust joint actuarial-financial framework for managing mortality risk in life insurance. They characterize the worst- and best-case contract values over all mortality processes that remain compatible with the observed life-table information and the financial market.
keywords:
Mortality modeling , Stochastic optimal control , Variable annuities , Life insurance , Risk management
[…]
1 Introduction
The two most common approaches to mortality modeling in life insurance rely either on fractional age assumptions or mortality rate models. Each provides a way to derive survival probabilities over continuous time, yet both come with inherent limitations.
(i) Fractional age assumptions. In practice, life insurers and pension funds often rely on life tables, which provide the survival probabilities of individuals at integer ages. However, many actuarial quantities (life annuities, variable annuities, or other functionals of the lifetime) depend on the full distribution of survival times, which requires information about the distribution of deaths between integer ages. To address this, actuaries introduce fractional age assumptions that interpolate survival probabilities between successive integer ages, see Chapter 3.6 of Bowers et al. (1997) and Chapter 3 of Dickson et al. (2020). The most common assumptions are the uniform distribution of deaths (UDD) assuming that deaths occur uniformly over each one-year interval; the constant force of mortality (CFM) with an exponential decay of the survival function within each interval; and the Balducci assumption under which the probability of death decreases linearly with time over each year. Each of these assumptions produces a different continuous-time survival distribution consistent with the same life table at integer ages. However, since life tables provide no information about the actual distribution of deaths between integer ages, the resulting actuarial quantities depend heavily on the chosen assumption, which introduces a structural source of model risk that is rarely quantified by insurers.
(ii) Mortality rate models. An alternative approach to fractional age assumptions is to model mortality continuously in time through a mortality rate, or force of mortality, $\mu_{x}(t)$ at initial age $x$ and to derive the corresponding survival function $S_{x}(t)=\exp\left(-\int_{0}^{t}\mu_{x}(s)ds\right)$. Deterministic mortality laws, such as the Gompertz–Makeham (Gompertz, 1825, Makeham, 1860) or Heligman–Pollard models (Heligman and Pollard, 1980), model $\mu_{x}(t)$ as a smooth function of time $t$ and provide analytical tractability for pricing and reserving purposes. More recent developments have introduced stochastic mortality models to capture systematic longevity trends and uncertainty in future mortality evolution. Notable examples include the Lee–Carter models (Lee and Carter, 1992, Booth et al., 2002, Renshaw and Haberman, 2003), the Cairns–Blake–Dowd (CBD) model (Cairns et al., 2006, 2009), and cohort-based extensions such as the Renshaw–Haberman model (Renshaw and Haberman, 2006, Currie, 2006) or multi-population approaches (Li and Lee, 2005), see also Plat (2009) and Basellini et al. (2023) for a thorough review on stochastic mortality modeling. These models specify continuous-time dynamics for mortality rates, enabling longevity risk management in a probabilistic framework. While both deterministic and stochastic mortality rate models define survival probabilities continuously for all $t\geq 0$, they are inherently model-based: their dynamics rely on parametric and structural assumptions about the underlying mortality process. Consequently, computed actuarial functionals of the lifetime depend strongly on these modeling choices, and different specifications can lead to substantially different valuations of life-contingent liabilities.
Therefore, to avoid these two sources of model incompleteness, authors have considered fractional-age interpolation and safe-side reserving — including the sum-at-risk tradition of Lidstone (1905), Norberg (1985), Linnemann (1993) and Kalashnikov and Norberg (2003); the fractional-age comparisons via stochastic ordering of Hürlimann (1990), Willmot (1997), Jones and Mereu (2000), Frostig (2002), Frostig (2003), Barz and Müller (2012); and the the worst-case techniques of Christiansen (2010), Christiansen and Denuit (2010, 2013). However, these approaches remain fundamentally actuarial in both object and method: it seeks to compare admissible lifetime distributions, construct conservative first-order mortality bases, or derive safe-side / worst-case actuarial reserves from one-year life-table information, typically through stochastic ordering, shape restrictions on mortality rates, Thiele-type reserve equations, and the sign of the sum at risk. Barz and Müller (2012) derive tight bounds for functionals of future lifetimes under an increasing force of mortality, while Christiansen and Denuit (2010) construct conservative first-order bases explicitly for safe-side actuarial calculations; Christiansen and Denuit (2013) extend this logic to single- and multiple-decrement settings and mixed survival/death benefits. What these papers do not provide is a joint financial-market framework: there are no traded assets, no self-financing investment strategy, no hedging argument, and no risk-neutral pricing measure. By contrast, the present paper keeps the same life-table incompleteness problem but shifts it into a market-consistent robust valuation setting, where mortality ambiguity is analyzed jointly with financial dynamics and contract pricing rather than only through actuarial reserve calculations. This, in turn, leads to the application of stochastic control techniques over mortality processes, which, to the best of our knowledge, is new to the literature and allows for a much broader admissible mortality class than the shape-restricted intensities of Barz and Müller (2012), Christiansen and Denuit (2010, 2013) and related work.
We therefore propose a framework that avoids both sources of model misspecification, namely, the arbitrary choice of fractional age assumption and the structural assumptions of mortality rate models, while being robust with modern financial valuation. We derive upper and lower bounds for actuarial functionals of the lifetime that are compatible with the observed life table at integer ages. Our approach characterizes all possible mortality rates consistent with these discrete data points, and determines the corresponding optimal values of hybrid actuarial-financial quantities such as variable annuities. We present two complementary formulations: a strict setting in which each mortality path must match almost surely all the one-year survival probabilities given by the reference life table, and a relaxed (more realistic) setting in which this constraint only holds in expectation, allowing for mortality trajectories to deviate while still being consistent on average with each one-year tabulated survival probabilities. More precisely, for a fixed maturity $T>0$ and a risk-neutral measure $\mathbb{Q}$, we aim at solving stochastic control problems of the form
$\sup_{\mu\in\mathcal{A}}\mathbb{E}^{\mathbb{Q}}\left[F(T,\mu,A,r)\right]\quad\ \text{and}\quad\ \inf_{\mu\in\mathcal{A}}\mathbb{E}^{\mathbb{Q}}\left[F(T,\mu,A,r)\right],$
for some functional $F$ of the lifetime [^1] Strictly speaking, functional of the mortality rates, but we interchangeably use these terms when it is clear from the context. $\mu=(\mu_{t})_{0\leq t\leq T}$, investment fund $A=(A_{t})_{0\leq t\leq T}$, and interest rates $r=(r_{t})_{0\leq t\leq T}$. The strict setting of Section 4 will define the set of admissible controls $\mathcal{A}$ such as to match almost surely the given survival probabilities over each integer year from the life table. As this strict setting effectively leads to deterministic optimal mortality rates, we then consider in Section 5 a relaxed setting where the admissible set $\mathcal{A}$ is such that this constraint only holds in expectation. However, the resulting control problems then tend to be ill-posed, and similar to Avellaneda and Paras (1996), Li and Szimayer (2011), we need to further restrict this set $\mathcal{A}$ to bounded controls from below and above by time-dependent functions consistent with the life table. We establish asymptotic results showing that, as the set of bounded controls expands, the solutions to the restricted problems converge to the solution of the unrestricted problems.
These two distinct yet complementary approaches let us derive general bounds for hybrid functionals of the lifetime that are compatible with observed life tables and with modern financial valuation, enabling insurers to quantify the worst- and best-case contract prices resulting from deviations of the observed mortality rates from their mortality assumptions/models. This framework is therefore not a pricing engine, but instead provides a novel robust and model-free risk management tool to quantify the sensitivity of financial life-contingent valuations to within-year mortality uncertainty. We highlight that ’model-free’ refers here to the mortality component: no explicit mortality model is ever imposed. A financial model is still required for valuation, see Section 3, but the bounds derived in Section 4 (strict setting) and 5 (relaxed setting) directly generalize to more complex financial dynamics. Numerical experiments in Section 6 further illustrate and support these findings.
2 Mortality modeling
Let $(\Omega,\mathcal{G},\mathbb{P})$ a probability space on which is defined the random variable $T_{0}:\Omega\to\mathbb{R}^{+}$ describing the future lifetime of a new born child. Let $x>0$, then $T_{x}:=\left(T_{0}-x\,\mid\,T_{0}>x\right):\Omega\to\mathbb{R}_{+}$ denotes the random variable describing the future lifetime of a person aged $x$. We consider in this paper a general framework where the lifetime is a mixed random variable with both discrete and continuous components. The cumulative distribution function (CDF) and survival function of $T_{x}$ are given by
$_{t}q_{x}:=\mathbb{P}[T_{x}\leq t]=F_{x}(t)\quad\text{and}\quad{}_{t}p_{x}:=\mathbb{P}[T_{x}>t]=S_{x}(t)\,.$
By definition, $F_{x}(t)$ is non-decreasing, right-continuous, and satisfies $F_{x}(0)=0$ with $\lim_{t\to\infty}F_{x}(t)=1$. Similarly, $S_{x}(t)$ is non-increasing, right-continuous, and satisfies $S_{x}(0)=1$ with $\lim_{t\to\infty}S_{x}(t)=0$. $_{t}p_{x}$ is then the survival probability of an $x$-year old individual till age $x+t$ , and $_{t}q_{x}$ his probability of dying before attaining time $t$. Furthermore, knowing the individual is still alive at age $x+t$, we write its survival probability to age $s+t$, $s\geq t$, as $_{s-t}p_{x+t}=\mathbb{P}[T_{x}>s+t\,|\,T_{x}>t]$. For an absolutely continuous random variable $T_{x}$, the probability density function (pdf) of $T_{x}$ is given by $f_{x}(t)=-dS_{x}(t)/dt$ and we define the force of mortality $\mu_{x,c}(t)$ as the conditional density function of $T_{x}$ at exact age $x+t$ given survival to that age,
$\mu_{x,c}(t)=f_{x}(t)/S_{x}(t)=-d\ln S_{x}(t)/dt\,$,
which is a continuous map $\mu_{x,c}:\mathbb{R}^{+}\to\mathbb{R}^{+}$. Clearly, we have
$_{t}p_{x}=\exp\left(-\int_{0}^{t}\mu_{x,c}(s)ds\right)\,\vskip 5.69054pt.$ (2.1)
On the other hand, if $T_{x}$ is a discrete random variable taking values $0<t_{0}<t_{1}<\ldots\in\mathbb{R}^{+}$ with associated probability function $f_{x}(t_{j})=\mathbb{P}(T_{x}=t_{j})$ for $j\in\mathbb{N}$, the probability of survival is
$_{t}p_{x}=\sum_{j\,:\,t_{j}>t}f_{x}(t_{j})\,.$
The (discrete) mortality rate $\mu_{x,d}:\mathbb{R}^{+}\to[0,1]$ is then defined accordingly as the conditional death probability at time $t_{j}$ given that the individual has survived to $t_{j}$,
$\mu_{x,d}(t_{j})=\mathbb{P}\left(T_{x}=t_{j}\,\mid\,T_{x}\geq t_{j}\right)=\frac{f_{x}(t_{j})}{S_{x}(t_{j}^{-})}\,$, (2.2)
where $S_{x}(t^{-})=\lim_{s\to t^{-}}S_{x}(s)$. When it is clear from the context, we simplify the notation of the discrete mortality process $\mu_{x,j}:=\mu_{x,d}(t_{j})$. Corresponding to (2.1), the survival probabilities can be written as
$_{t}p_{x}=\prod_{j\,:\,t_{j}\leq t}(1-\mu_{x,j})\quad\text{and}\quad f_{x}(t_{j})=\mu_{x,j}\prod_{i=1}^{j-1}(1-\mu_{x,i})\,.$ (2.3)
As in the continuous case, the discrete hazard process $(\mu_{x,j})_{j\in\mathbb{N}}$ uniquely determines the distribution of the lifetime $T_{x}$.
More generally, the distribution of $T_{x}$ may have both discrete and continuous components. In this case, the mortality rate can be defined to have continuous component $\mu_{x,c}(t)$ and discrete components $\mu_{x,0},\mu_{x,1},\ldots$ at the discrete times $t_{0},t_{1},\ldots$. From Kalbfleisch and Prentice (2002), the survival probability can then be rewritten
$_{t}p_{x}=\exp\left(-\int_{0}^{t}\mu_{x,c}(s)ds\right)\prod_{j\,:\,t_{j}\leq t}(1-\mu_{x,j})\,$, (2.4)
and as before, the conditional probability of death over $[t,t+dt)$ is given by the hazard measure:
$\mathbb{P}[T_{x}\in[t,t+dt)\mid T_{x}\geq t]=\begin{cases}\mu_{x,j},&t=t_{j},\quad j=0,1,2,\ldots\\ \mu_{x,c}(t)\,dt,&\text{ otherwise.}\end{cases}\vskip 5.69054pt$ (2.5)
Instead of the deterministic setting (2.4), we allow in this work for stochastic mortality intensity. Specifically, we consider a continuous-time component $\mu_{x,c}=(\mu_{x,c}(t))_{0\leq t\leq T}$ and a discrete-time component $\mu_{x,d}=(\mu_{x,d}(\tau_{j}))_{j\in\mathbb{N}}$. Both processes are assumed to be adapted to the filtration $\mathbb{G}:=(\mathcal{G}_{t})_{0\leq t\leq T}$ such that $\mathcal{G}_{T}\subset\mathcal{G}$, with $(\tau_{j})_{j\in\mathbb{N}}$ a strictly increasing sequence of $\mathbb{G}$-stopping time with $0<\tau_{0}<\tau_{1}<\cdots$ a.s. (replacing the deterministic time $t_{j}$) and the jump size $\mu_{x,j}=\mu_{x,d}(\tau_{j})\in[0,1]$ being $\mathcal{G}_{\tau_{j}}$-measurable for each $j$. We then extend (2.4) using
$_{t}p_{x}=\mathbb{E}\hskip-1.42262pt\left[\exp\left(-\int_{0}^{t}\mu_{x,c}(s)ds\right)\hskip-1.42262pt\prod_{j\,:\,\tau_{j}\leq t}(1-\mu_{x,j})\ \Big|\ \mathcal{G}_{0}\right],$ (2.6)
and the probability of survival till time $T$ being alive at time $t+x$,
$_{T-t}p_{x+t}=\mathbb{E}\hskip-1.42262pt\left[\exp\left(-\int_{t}^{T}\mu_{x,c}(s)ds\right)\hskip-1.42262pt\prod_{j\,:\,t<\tau_{j}\leq T}(1-\mu_{x,j})\ \Big|\ \mathcal{G}_{t}\right].$ (2.7)
Finally, we assume that we are given some life table which is described as follows. Let us have a group of $l_{0}$ new born children. Then, for $\mathds{1}_{i}(x)$ indicating the survival of the new born child number $i$ to age $x$, $L(x):=\sum_{i=1}^{l_{0}}\mathds{1}_{i}(x)$ is a random variable denoting the number of children alive at age $x$. For every $i=1,\ldots,l_{0}$ and with $_{x}p_{0}:=\mathbb{P}[\mathds{1}_{i}(x)=1]$, the expectation of $L(x)$ is
$l_{x}:=\mathbb{E}[L(x)]=\mathbb{E}\left[\sum_{i=1}^{l_{0}}\mathds{1}_{i}(x)\right]=l_{0}\cdot{}_{x}p_{0}\,$,
which is precisely the quantity reported in the life table. More generally, we have for any integer $j\in\mathbb{N}$,
$l_{x+j}=l_{0}\cdot{}_{x+j}p_{0}=l_{0}\cdot{}_{x}p_{0}\cdot{}_{j}p_{x}=l_{x}\cdot{}_{j}p_{x}\,.$ (2.8)
Therefore, we can equivalently consider that the given life table provides us with $_{j}p_{x}=l_{x+j}/l_{x}$ for $j\in\mathbb{N}$, and thus only specifies the survival function of $\lfloor T(x)\rfloor$. However, many actuarial products are functionals of the lifetime, and require full information on the the survival function of $T(x)$, i.e. on $_{t}p_{x}$ for all $t\in\mathbb{R}^{+}$. The insurer must then postulate an analytic form for the mortality rates (Gompertz-Makeham, Lee-Carter, CBD, etc.) or adopt a fractional age assumption in addition to the given life tables. We refer to Chapters 2 and 3 of Dickson et al. (2020) for a thorough review on mortality modeling, and give here a brief reminder of the three main fractional age assumptions. Given a fixed $x=0,1,\ldots$ and [^2] We restrict $t\in[0,1)$ since w.l.o.g. when $t>1$, we can always rewrite $_{t}p_{x}={}_{\lfloor t\rfloor}p_{x}\cdot{}_{t^{\prime}}p_{x+\lfloor t\rfloor}$ with $t^{\prime}=t-\lfloor t\rfloor\in(0,1)$. $t\in[0,1)$, we have
  • Uniform distribution of deaths (UDD),
    $_{t}p_{x}=1-t\cdot(1-{}_{1}p_{x})\,.$ (2.9)
  • Constant force of mortality (CFM),
    $_{t}p_{x}=e^{-t\cdot\mu_{x}(0)}=({}_{1}p_{x})^{t}\,.$ (2.10)
  • Balducci approximation,
    $_{t}p_{x}=\frac{{}_{1}p_{x}}{t+(1-t)\cdot{}_{1}p_{x}}\,.$ (2.11)
3 The financial market
We then consider the following financial framework for illustrating purposes, while emphasizing that the contract bounds derived in Section 4 and 5 can be extended directly to other and more complex (Markovian) financial dynamics [^3] This is straightforward by adapting the contract price $C_{t}$ in Section 4 and the coefficients of the HJB equations in Section 5.. The interest rate and stock processes, $\left(r_{t}\right)_{0\leq t\leq T}$ and $\left(S_{t}\right)_{0\leq t\leq T}$ are defined on $(\Omega,\mathcal{F},\mathbb{F}:=(\mathcal{F}_{t})_{0\leq t\leq T},\mathbb{Q})$ generated by $\mathbf{W}^{\mathbb{Q}}=\left(W_{t}^{1,\mathbb{Q}},W_{t}^{2,\mathbb{Q}}\right)_{0\leq t\leq T}^{\top}$ :
$d\,\binom{\,S_{t}\,}{r_{t}}=\binom{r_{t}S_{t}}{\ \kappa\left(\theta(t)-r_{t}\right)\ }\,dt+\binom{\ S_{t}\,\sigma_{S}\,\mathbf{e}_{1}^{\top}\ }{\sigma_{r}\,\mathbf{e}_{2}^{\top}}\,\Sigma\,d\mathbf{W}^{\mathbb{Q}}_{t}\,$, (3.1)
where $\mathbf{e}_{1}^{\top}=(1,0),\mathbf{e}_{2}^{\top}=(0,1).\ \Sigma$ is the Cholesky decomposition of the correlation matrix between the stock and the interest rate,
$\Sigma=\left(\begin{array}[]{cc}1&0\\ \rho&\sqrt{1-\rho^{2}}\end{array}\right),\ \quad\Sigma\Sigma^{\top}=\left(\begin{array}[]{cc}1&\rho\\ \rho&1\end{array}\right),$
where $\rho\in(-1,1)$ is the stock-rate correlation. The cash account is $B_{t}=\exp\left(\int_{0}^{t}r_{s}ds\right)$ and the zero-coupon (ZC) bond $P(t,T)=\mathbb{E}^{\mathbb{Q}}\left[e^{-\int_{t}^{T}r_{s}ds}\,\big|\,\mathcal{F}_{t}\right]$. Then this financial market is complete and we know that there exists a unique density process $(Z_{t})_{0\leq t\leq T}$ defined as
$Z_{t}:=\frac{d\mathbb{Q}}{d\mathbb{P}}\bigg|_{\mathcal{F}_{t}},\quad 0\leq t\leq T.$ (3.2)
Note that the mortality rate processes $\mu_{x,c},\mu_{x,d}$ (and thus the random future lifetime $T_{x}$) are not assumed independent from the financial assets $\left(r_{t}\right)_{0\leq t\leq T}$ and $\left(S_{t}\right)_{0\leq t\leq T}$, see Assumption 4.2. The parameters are $\kappa,\sigma_{S},\sigma_{r}>0$ and $\theta(t)$ is a function of time, chosen to match the yield curve
$\theta(t)=\frac{1}{\kappa}\frac{\partial}{\partial t}f(0,t)+f(0,t)+\frac{\sigma_{r}^{2}}{2\kappa^{2}}\left(1-e^{-2\kappa t}\right),$ (3.3)
where $f(0,t)=-\partial_{t}\ln P(0,t)$ is the instantaneous forward rate. With this specification, the interest rate model can be fitted on a standalone basis. The next standard proposition, see e.g. Hainaut (2024), reminds the price of a ZC bond in this model.
Proposition 3.1.
The price at time $t$ of a zero-coupon bond of maturity $T\geq t$, is given by
$P(t,T)=\exp\left(-B(t,T)r_{t}+\ln\frac{P(0,T)}{P(0,t)}+B(t,T)f(0,t)-\frac{\sigma_{r}^{2}}{4\kappa^{3}}\left(1-e^{-2\kappa t}\right)B(t,T)^{2}\right),$
where $B(t,T)=\frac{1}{\kappa}\left(1-e^{-\kappa(T-t)}\right)$. Consequently, under the risk-neutral measure $\mathbb{Q}$, we obtain the bond price dynamics as
$\frac{dP(t,T)}{P(t,T)}=r_{t}dt-B(t,T)\,\sigma_{r}\,\mathbf{e}_{2}^{\top}\Sigma\,d\mathbf{W}^{\mathbb{Q}}_{t}\,.$ (3.4)
The policyholder’s premium is then invested in a fund $(A_{t})_{0\leq t\leq T}$ composed of cash, stocks and ZC bonds. More precisely, the insurer invests constant percentages $\pi_{S}$ in stocks, $\pi_{P}$ in bonds and $1-\pi_{S}-\pi_{P}$ in cash ($\bm{\pi}=(\pi_{S},\pi_{P})\in[0,1]^{2}$ and $\left.\pi_{S}+\pi_{P}\leq 1\right)$. The fund is self-financed (no exogenous infusion or withdrawal of money) and its instantaneous return is given by
$\frac{dA_{t}}{A_{t}}=\pi_{S}\frac{dS_{t}}{S_{t}}+\pi_{P}\frac{dP(t,T)}{P(t,T)}+(1-\pi_{S}-\pi_{P})\,r_{t}dt\,.$ (3.5)
If we remember Equation (3.1) and (3.4), we infer that
$\frac{dA_{t}}{A_{t}}$ $=r_{t}dt+\pi_{S}\sigma_{S}\mathbf{e}_{1}^{\top}\Sigma\,d\mathbf{W}^{\mathbb{Q}}_{t}-\pi_{P}\sigma_{r}B(t,T)\,\mathbf{e}_{2}^{\top}\Sigma\,d\mathbf{W}^{\mathbb{Q}}_{t}$ (3.6)
$=r_{t}dt+\left(\pi_{S}\sigma_{S}\mathbf{e}_{1}^{\top}-\pi_{P}\sigma_{r}B(t,T)\,\mathbf{e}_{2}^{\top}\right)\Sigma\,d\mathbf{W}^{\mathbb{Q}}_{t}\,.$
Let us consider for now the pricing of a financial derivative with payoff
$H(A_{T},G_{T})=\max\{A_{T},G_{T}\}=G_{T}+(A_{T}-G_{T})_{+}\,.$ (3.7)
A typical choice in the following consists in $G_{T}=A_{0}e^{r_{g}T}$ for $r_{g}\geq 0$ a minimum guaranteed return. The price of this contract with maturity $T$ is equal to the expected discounted cash-flow under the risk-neutral measure $\mathbb{Q}$,
$C_{t}(T,G_{T})=\mathbb{E}^{\mathbb{Q}}\left[e^{-\int_{t}^{T}r_{s}ds}\,H(A_{T},G_{T})\,\big|\,\mathcal{F}_{t}\right].$ (3.8)
The next Proposition follows from Hainaut (2024) and gives us the price of this contract.
Proposition 3.2.
The contract’s price $C_{t}$ for $t\leq T$ is given by
$C_{t}(T,G_{T})=G_{T}P(t,T)+A_{t}\Phi(-d_{1}(t))-G_{T}P(t,T)\Phi(-d_{2}(t))\,$, (3.9)
where
$d_{1}(t)$ $=d_{2}(t)-\sigma_{Y}(t,T)\ ,\quad\ \ d_{2}(t)=\frac{\ln(\frac{G_{T}P(t,T)}{A_{t}})+\frac{1}{2}\sigma^{2}_{Y}(t,T)}{\sigma_{Y}(t,T)}\,$, (3.10)
$\sigma^{2}_{Y}(t,T)$ $=\int_{t}^{T}\left(B(s,T)\sigma_{r}\mathbf{e}_{2}^{\top}\Sigma+\bm{\pi}^{\top}\bm{\Sigma}_{F}(s)\right)\left(B(s,T)\sigma_{r}\Sigma^{\top}\mathbf{e}_{2}+\bm{\Sigma}_{F}^{\top}(s)\bm{\pi}\right)ds\,$, (3.11)
$\bm{\Sigma}_{F}(s)$ $=\left(\begin{array}[]{cc}\sigma_{S}&0\\ -\rho\sigma_{r}B(s,T)&-\sqrt{1-\rho^{2}}\,\sigma_{r}B\left(s,T\right)\end{array}\right).$ (3.14)
Note that (3.11) can be obtained in closed-form, see Appendix B and C of Hainaut (2024).
4 Bounds on variable annuities: almost sure matching of the life tables
In this section, we first derive bounds for three variable annuity products with respect to mortality rates, while enforcing strict consistency with the observed life tables. Accordingly, we assume that the total force of mortality available for each integer year $x$ (or equivalently, each one-year survival probability) is given by the life table $(l_{x})$ in (2.8) and we then restrict attention to admissible mortality trajectories that reproduce almost surely each of these prescribed annual quantities, see constraint (4.6).
Let us assume that we are given a cohort of individuals aged $x$ at time $t=0$ and the augmented filtration $\mathbb{H}=(\mathcal{H}_{t})_{0\leq t\leq T}$ given by $\mathcal{H}_{t}:=\mathcal{F}_{t}\vee\mathcal{G}_{t}$. We require the two following standard assumptions throughout the rest of the manuscript.
Assumption 4.1.
The random variable $T_{x}$ satisfies for $t\leq s\leq T,$
$\mathbb{P}(T_{x}>s\mid\mathcal{H}_{s}\vee\{T_{x}>t\})=e^{-\int_{t}^{s}\mu_{x,c}(u)du}\hskip-5.69054pt\prod_{j\,:\,t<\tau_{j}\leq s}\hskip-5.69054pt(1-\mu_{x,j})\,.$ (4.1)
This property is an extension of the doubly stochastic (Cox) property, see Cox and Isham (1980).
Assumption 4.2.
In the combined financial-actuarial market, we assume that $\mathbb{Q}$ corresponds to the Minimal Martingale Measure (MMM) (Schweizer, 1991, 1995). Under this assumption, the density of the change of measure is such that
$\frac{d\mathbb{Q}}{d\mathbb{P}}\bigg|_{\mathcal{H}_{t}}=Z_{t},\quad 0\leq t\leq T,$ (4.2)
where $(Z_{t})_{0\leq t\leq T}$ is the unique density process defined by (3.2).
Several arguments support the use of the MMM. In particular, in the valuation and hedging of variable annuities, Møller (1998, 2001) and Ceci et al. (2017) show that, within the framework of local risk-minimization, the fair value of variables annuities can be expressed as the discounted expectation of its payoff under the MMM. Hence, this measure arises naturally in incomplete markets, where perfect replication is not feasible, and it provides the pricing rule consistent with the locally risk-minimizing hedging strategy. The MMM is also mathematically tractable: its construction relies solely on the orthogonal decomposition of the discounted price process. Moreover, under the MMM, no additional mortality risk premium is introduced. More precisely, since the mortality intensity may be correlated with traded assets, its drift is adjusted under the MMM, which induces a change in the dynamic but not in the functional form of the mortality intensity. Although the choice of the martingale measure in Assumption 4.2 appears arbitrary, it is in fact the most natural one in our setting as we have no prior knowledge regarding the mortality rates $\mu$. Other martingale measures could be considered but would require additional modeling choices, such as the mortality risk process $(\lambda_{t})_{0\leq t\leq T}$, with $\lambda_{t}>-1$ a.s., which determines the pricing mortality under $\mathbb{Q}$ as $\mu^{\mathbb{Q}}_{t}=(1+\lambda_{t})\mu_{t}$ for all $0\leq t\leq T$, as well as assumptions on any other stochastic factors that could influence the dynamics of mortality process.
Assumption 4.2 justifies to introduce our general objective functional as the expectation under $\mathbb{Q}$ of a given functional $F$ of the lifetime (equivalently, of the mortality rates), account value $(A_{t})_{0\leq t\leq T}$ and interest rate $(r_{t})_{0\leq t\leq T}$, as follows
$V^{\mu}_{t}=\mathbb{E}^{\mathbb{Q}}\left[F(T,\mu_{x,c},\mu_{x,d},A,r)\mid\mathcal{H}_{t}\right],\qquad t\in[0,T].$ (4.3)
Our objective is to derive lower and upper bounds for $V^{\mu}_{t}$ over a prescribed class of mortality models consistent with a given life table $(l_{x})_{x\in\mathbb{N}}$ described by (2.8). To this end, we consider the continuous-time force of mortality $\mu_{x,c}=(\mu_{x,c}(t))_{0\leq t\leq T}$ as a control process taking values in a given compact set $\mathbb{M}\subset[0,+\infty)$, and we let $\mathcal{M}$ be a given family of such continuous $\mathbb{H}$-adapted controls $\mu_{x,c}$. In addition, we consider the discrete process $\mu_{x,d}=(\tau_{0},\tau_{1},\ldots;\mu_{x,0},\mu_{x,1},\ldots)\in\mathcal{V}$ as a given impulse control, where $\mathcal{V}$ is the set of admissible impulse controls $\mu_{x,d}$ such that $0<\tau_{0}<\tau_{1}<\ldots\leq T$ (a.s.) are $\mathbb{H}$-stopping times and $\mu_{x,j}\in[0,1]$ are $\mathcal{H}_{\tau_{j}}$-measurable for all $j\geq 0$. We then call $\mu:=(\mu_{x,c},\mu_{x,d})$ a combined control. Finally, we assume that we are given a set $\mathcal{W}\subset\mathcal{M}\times\mathcal{V}$ of admissible combined controls $\mu$, such that $V^{\mu}_{t}$ is well-defined for all $t\in[0,T]$ and such that the life-table constraint (4.6) below is satisfied.
Note that we do not only want the force of mortality to be consistent with the final survival probability $_{T-t}p_{x+t}$, but with all intermediate one-year survival probabilities at integer ages between $x+t$ and $x+T$. This can be seen writing for $t,T\in\mathbb{R}^{+}$,
$_{T-t}p_{x+t}={}_{\lceil t\rceil-t}p_{x+t}\times{}_{1}p_{x+\lceil t\rceil}\times{}_{1}p_{x+1+\lceil t\rceil}\times\cdots\times{}_{1}p_{x+j+\lceil t\rceil}\times\cdots\times{}_{T-\lfloor T\rfloor}p_{x+\lfloor T\rfloor}\,.$
All the one-year survival probabilities $_{1}p_{x+j+\lceil t\rceil}$, $j=0,\ldots,\lfloor T\rfloor-\lceil t\rceil-1$, can be observed from the given life table (2.8), i.e.
$\hat{p}_{j}:={}_{1}p_{x+j+\lceil t\rceil}=\frac{l_{x+j+\lceil t\rceil+1}}{l_{x+j+\lceil t\rceil}}.$ (4.4)
Hence, we obtain [^4] Additional constraints from the observed life table (such as fractional age assumptions or from expert knowledge) can also be imposed on $_{\lceil t\rceil-t}p_{x+t}$ and $_{T-\lfloor T\rfloor}p_{x+\lfloor T\rfloor}$. Alternatively, we simply assume in the following that these extremal survival probabilities are fixed to $1$ and cannot be controlled in our framework through $\mu$.
$_{T-t}p_{x+t}={}_{\lceil t\rceil-t}p_{x+t}\times\hat{p}_{0}\times\hat{p}_{1}\times\cdots\times\hat{p}_{j}\times\cdots\times{}_{T-\lfloor T\rfloor}p_{x+\lfloor T\rfloor}.$ (4.5)
Therefore, in the strict setting of this section, admissible mortality controls $\mu\in\mathcal{W}$ are required to strictly match these observed $\hat{p}_{j}$’s. Specifically, for each $j=0,\ldots,\lfloor T\rfloor-\lceil t\rceil-1$,
$\exp\left(-\int_{\lceil t\rceil+j}^{\lceil t\rceil+j+1}\mu_{x,c}(s)ds\right)\hskip-3.69885pt\prod_{i\,:\,\lceil t\rceil+j<\tau_{i}\leq\lceil t\rceil+j+1}\hskip-11.38109pt(1-\mu_{x,i})=\hat{p}_{j}\,,\quad\mathbb{P}-\text{a.s.}$ (4.6)
The aim of this Section 4 is then to find the upper bound
$\bar{V}_{t}=\sup_{\mu\in\mathcal{W}}\mathbb{E}^{\mathbb{Q}}\left[F(\mu_{x,c},\mu_{x,d},A,r)\mid\mathcal{H}_{t}\right],$ (4.7)
and the lower bound
$\underline{V\mkern-4.0mu}\mkern 4.0mu_{t}=\inf_{\mu\in\mathcal{W}}\mathbb{E}^{\mathbb{Q}}\left[F(\mu_{x,c},\mu_{x,d},A,r)\mid\mathcal{H}_{t}\right],$ (4.8)
of functionals of the lifetime over all admissible mortality controls. We then denote, if they exist, the corresponding optimal mortality controls $\bar{\mu}^{*},\underline{\mu\mkern-4.0mu}\mkern 4.0mu^{*}\in\mathcal{W}$ attaining these bounds, i.e. $\bar{V}_{t}=V^{\bar{\mu}^{*}}_{t}$ and $\underline{V\mkern-4.0mu}\mkern 4.0mu_{t}=V^{\underline{\mu\mkern-4.0mu}\mkern 4.0mu^{*}}_{t}$.
Moreover, we emphasize that this pathwise constraint is quite strong in practice: it requires almost surely that every realized mortality path matches the annual survival probabilities $\hat{p}_{j}$ from the life table, essentially leading to deterministic optimal rates $\bar{\mu}^{*},\underline{\mu\mkern-4.0mu}\mkern 4.0mu^{*}$ (see Propositions 4.34.224.5 below). While a constraint in expectation would be more natural, a naive implementation of the resulting control problem appears to be ill-posed. We address this issue in Section 5, by appropriately relaxing this constraint (4.6) and specifying the corresponding set of admissible (bounded) controls. In the following subsections, we study the upper and lower bounds of three different types of variables annuities: GMAB, GMIB and GMDB. Proofs are deferred in A.
4.1. Guaranteed Minimum Accumulation Benefits
Guaranteed Minimum Accumulation Benefits (GMAB) are retirement savings products which promises in case of survival, the maximum between investments and a guaranteed capital. The contract is subscribed by an individual aged $x$ and guarantees at expiry date $T$ a payout equal to the maximum between the guaranteed capital $G^{A}_{T}=A_{0}e^{r_{g}T}$ and the fund $A_{T}$, in the event of survival. The payout in case of survival is then given by (3.7). Assuming survival to time tt [^5] We slightly abuse notation throughout the manuscript by conditioning on $\mathcal{H}_{t}$ instead of $\mathcal{H}_{t}\vee\{T_{x}>t\}$., the fair value of such a policy at time $t$ is denoted by $V^{\mu}_{t}$ and equal to the expected discounted cash-flows under $\mathbb{Q}$,
$V^{\mu}_{t}$ $=\mathbb{E}^{\mathbb{Q}}\left[e^{-\int_{t}^{T}r_{s}ds}\,\mathds{1}_{\{T_{x}>T\}}\,H(A_{T},G^{A}_{T})\ \big|\ \mathcal{H}_{t}\right]$
$=\mathbb{E}^{\mathbb{Q}}\left[\mathbb{E}^{\mathbb{Q}}\left[e^{-\int_{t}^{T}r_{s}ds}\,\mathds{1}_{\{T_{x}>T\}}\,H(A_{T},G^{A}_{T})\ \big|\ \mathcal{H}_{T}\right]\,\Big|\ \mathcal{H}_{t}\right]$
$=\mathbb{E}^{\mathbb{Q}}\bigg[\,e^{-\int_{t}^{T}(r_{s}+\mu_{x,c}(s))ds}\hskip-5.69054pt\prod_{j\,:\,t<\tau_{j}\leq T}\hskip-5.69054pt(1-\mu_{x,j})\,H(A_{T},G^{A}_{T})\ \big|\ \mathcal{H}_{t}\,\bigg],$ (4.9)
using Assumptions 4.1 and 4.2. Note that in general $V^{\mu}_{t}\neq\mathbb{E}^{\mathbb{Q}}[\mathds{1}_{T_{x}>T}\,|\,\mathcal{G}_{t}]\,C_{t}(T,G^{A}_{T})={}_{T-t}p_{x+t}\,C_{t}(T,G^{A}_{T})$ since $\mu$ is not assumed independent from $A$ and $r$. We then derive the GMAB upper bound
$\bar{V}_{t}=\sup_{\mu\in\mathcal{W}}\mathbb{E}^{\mathbb{Q}}\left[e^{-\int_{t}^{T}(r_{s}+\mu_{x,c}(s))ds}\hskip-5.69054pt\prod_{j\,:\,t<\tau_{j}\leq T}\hskip-5.69054pt(1-\mu_{x,j})\,H(A_{T},G^{A}_{T})\ \big|\ \mathcal{H}_{t}\right],$ (4.10)
and GMAB lower bound
$\underline{V\mkern-4.0mu}\mkern 4.0mu_{t}=\inf_{\mu\in\mathcal{W}}\mathbb{E}^{\mathbb{Q}}\left[e^{-\int_{t}^{T}(r_{s}+\mu_{x,c}(s))ds}\hskip-5.69054pt\prod_{j\,:\,t<\tau_{j}\leq T}\hskip-5.69054pt(1-\mu_{x,j})\,H(A_{T},G^{A}_{T})\ \big|\ \mathcal{H}_{t}\right].$ (4.11)
However, these GMAB bounds are trivial as any force of mortality $\mu\in\mathcal{W}$–thereby satisfying (4.6)–leads to the same price $V^{\mu}_{t}=V_{t}$, since the contract’s value only depends on the total survival probability $_{T-t}p_{x+t}$ given by (4.5). In fact, the exact death time of an individual in each one-year period does not change the contract’s value $V^{\mu}_{t}$ as in any case, a deceased policyholder will not receive the terminal payoff $H(A_{T},G^{A}_{T})$. In other words, the GMAB does not depend on the exact path of the mortality process, as summarized in the following proposition.
Proposition 4.3 (GMAB bounds).
For the GMAB contract (4.9), each admissible control $\mu\in\mathcal{W}$ is an optimal control $\bar{\mu}^{*}$ of (4.10) and $\underline{\mu\mkern-4.0mu}\mkern 4.0mu^{*}$ of (4.11). The GMAB price then satisfies for each $\mu\in\mathcal{W}$,
$\bar{V}_{t}=V^{\mu}_{t}={}_{T-t}p_{x+t}\,C_{t}(T,G^{A}_{T})\,$, (4.12)
and, respectively,
$\underline{V\mkern-4.0mu}\mkern 4.0mu_{t}=V^{\mu}_{t}={}_{T-t}p_{x+t}\,C_{t}(T,G^{A}_{T})\,.$ (4.13)
where $_{T-t}p_{x+t}$ is given by (4.5).
This Proposition 4.3 is still valid when considering any form of payoff $H(A_{T},G^{A}_{T})$ for the GMAB (4.9), by adjusting the price $C_{t}(T,G^{A}_{T})$ in (4.12)-(4.13). Even though the bounds on the GMAB are trivial, they lay the foundations for studying the following more complex variable annuities.
4.2. Guaranteed Minimum Income Benefit
We now consider the following variant of Guaranteed Minimum Income Benefits (GMIB) where an individual aged $x$ continuously receives a guaranteed rate $r_{g}>0$ on the initial investment $A_{0}$ as long as he is alive, plus a bonus rate proportional to the excess return of the fund $A_{t}$ over the pre-specified guarantee $G^{I}_{t}=A_{0}e^{r_{g}t}$. The exact payoff is given by666one could also consider continuously-compounded rates with $\widetilde{H}(A_{t},G^{I}_{t})=A_{0}\left(r_{g}e^{r_{g}(T-t)}+\frac{1}{A_{0}}(A_{t}-G^{I}_{t})_{+}\right)\,$ since $\int_{t}^{T}r_{g}e^{r_{g}(T-s)}ds=e^{r_{g}(T-t)}-1$.
$\widetilde{H}(A_{t},G^{I}_{t})=A_{0}\left(r_{g}+\frac{1}{A_{0}}(A_{t}-G^{I}_{t})_{+}\right)\,.$ (4.14)
Proposition 3.2 can be straightforwardly adapted to obtain as contract price for $s\geq t$, the sum of a ZC bond and a European call,
$C_{t}(s,G^{I}_{s})$ $=\mathbb{E}^{\mathbb{Q}}\left[e^{-\int_{t}^{s}r_{u}du}\,\widetilde{H}(A_{s},G^{I}_{s})\,\big|\,\mathcal{F}_{t}\right]$
$=A_{0}r_{g}P(t,s)+A_{t}\Phi(-d_{1}(t))-G^{I}_{s}P(t,s)\Phi(-d_{2}(t)).$ (4.15)
Again, the exact form of the payoff $\widetilde{H}$ is chosen here solely for analytical convenience in solving the optimization (4.17)–(4.18) and does not restrict the generality of our analysis as Proposition 4.22 below remains valid for any positive payoff $\widetilde{H}$, provided that the corresponding price $C_{t}(s,G^{I}_{s})$ is adjusted accordingly. The fair value of such a policy, denoted by $V^{\mu}_{t}$, is again equal to the expected discounted cash-flows under the risk neutral measure $\mathbb{Q}$,
$V^{\mu}_{t}$ $=\mathbb{E}^{\mathbb{Q}}\left[\int_{t}^{T}\hskip-2.84526pte^{-\int_{t}^{s}r_{u}du}\,\mathds{1}_{T_{x}>s}\,\widetilde{H}(A_{s},G^{I}_{s})\,ds\ \Big|\ \mathcal{H}_{t}\right]\hskip-1.99168pt$
$=\mathbb{E}^{\mathbb{Q}}\left[\int_{t}^{T}\hskip-2.84526pte^{-\int_{t}^{s}r_{u}du}\,{}_{s-t}p_{x+t}\,\widetilde{H}(A_{s},G^{I}_{s})\,ds\ \Big|\ \mathcal{H}_{t}\right]$
$=\mathbb{E}^{\mathbb{Q}}\left[\int_{t}^{T}e^{-\int_{t}^{s}(r_{u}+\mu_{x,c}(u))du}\hskip-5.69054pt\prod_{j\,:\,t<\tau_{j}\leq s}\hskip-5.69054pt(1-\mu_{x,j})\,\widetilde{H}(A_{s},G^{I}_{s})\,ds\ \Big|\ \mathcal{H}_{t}\right],$ (4.16)
under Assumptions 4.1 and 4.2. The upper bound
$\bar{V}_{t}=\sup_{\mu\in\mathcal{W}}\mathbb{E}^{\mathbb{Q}}\left[\int_{t}^{T}e^{-\int_{t}^{s}(r_{u}+\mu_{x,c}(u))du}\hskip-5.69054pt\prod_{j\,:\,t<\tau_{j}\leq s}\hskip-5.69054pt(1-\mu_{x,j})\,\widetilde{H}(A_{s},G^{I}_{s})\,ds\ \Big|\ \mathcal{H}_{t}\right],$ (4.17)
and lower bound
$\underline{V\mkern-4.0mu}\mkern 4.0mu_{t}=\inf_{\mu\in\mathcal{W}}\mathbb{E}^{\mathbb{Q}}\left[\int_{t}^{T}e^{-\int_{t}^{s}(r_{u}+\mu_{x,c}(u))du}\hskip-5.69054pt\prod_{j\,:\,t<\tau_{j}\leq s}\hskip-5.69054pt(1-\mu_{x,j})\,\widetilde{H}(A_{s},G^{I}_{s})\,ds\ \Big|\ \mathcal{H}_{t}\right],$ (4.18)
of the GMIB contract (4.16) are again such that the constraint (4.6) from the given life table holds and are derived in the following proposition.
Proposition 4.4 (GMIB bounds).
(i) Upper bound. The optimal control $\bar{\mu}^{*}=(\bar{\mu}^{*}_{x,c},\bar{\mu}^{*}_{x,d})$ for the upper bound $\bar{V}_{t}$ is given by
$\bar{\mu}^{*}_{x,c}(t)\equiv 0,\qquad\bar{\mu}^{*}_{x,d}=\{(\bar{\tau}_{j}^{*},\bar{\mu}_{x,j}^{*})\}_{j=0,\dots,\lfloor T\rfloor-\lceil t\rceil-1},$
with intervention times and jump sizes
$\bar{\tau}_{j}^{*}=\lceil t\rceil+j+1,\qquad\bar{\mu}_{x,j}^{*}=1-\hat{p}_{j}.$
Moreover, the upper bound satisfies
$\bar{V}_{t}=\int_{t}^{T}{}_{s-t}\bar{p}^{*}_{x+t}\,C_{t}(s,G_{s}^{I})\,ds,$ (4.19)
where the induced survival probability is deterministic and is equal to
$_{s-t}\bar{p}^{*}_{x+t}=\prod_{j\in\mathbb{N}:\,\lceil t\rceil+j+1\leq s}\hat{p}_{j},$ (4.20)
for $s\in[t,T]$ (with the convention $\prod_{\emptyset}:=1$).
(ii) Lower bound. The lower bound $\underline{V}_{t}:=\inf_{\mu\in\mathcal{W}}V_{t}^{\mu}$ is not attained in $\mathcal{W}$. However, there exists a sequence of admissible controls $\{\mu^{(n)}\}_{n\geq 1}\subset\mathcal{W}$ such that $V_{t}^{\mu^{(n)}}\downarrow\underline{V}_{t}$ and whose discrete intervention times satisfy $\tau^{(n)}_{j}\downarrow\lceil t\rceil+j$ with $\mu^{(n)}_{x,j}=1-\hat{p}_{j}$ for each $j=0,\dots,\lfloor T\rfloor-\lceil t\rceil-1$ and with $\mu_{x,c}^{(n)}\equiv 0$. This lower bound $\underline{V\mkern-4.0mu}\mkern 4.0mu_{t}$ then satisfies
$\underline{V\mkern-4.0mu}\mkern 4.0mu_{t}=\int_{t}^{T}{}_{s-t}\underline{p\mkern-4.0mu}\mkern 4.0mu^{*}_{x+t}\,C_{t}(s,G^{I}_{s})\,ds\,$, (4.21)
where for $s\in[t,T]$,
${}_{s-t}\underline{p\mkern-4.0mu}\mkern 4.0mu^{*}_{x+t}=\prod_{j\in\mathbb{N}\,:\,\lceil t\rceil+j<s}\hat{p}_{j}\,.$ (4.22)
We can extend Proposition 4.22 to more general payoffs by allowing $\widetilde{H}$ to be negative, e.g. if $r_{g}\in\mathbb{R}$. The existence and uniqueness of each $\tau^{*}_{j}$ then depend on the growth and convexity properties of the $\mathcal{H}_{t}$-conditional expectation in (4.16). Naturally, when $\tau^{*}_{j}$ is unique, we obtain directly from constraint (4.6) with $\mu^{*}_{x,c}\equiv 0$ that $\mu^{*}_{x,j}=1-\hat{p}_{j}$. Even when $\tau^{*}_{j}$ is not unique, choosing arbitrarily any such $\tau^{*}_{j}$ with $\mu^{*}_{x,j}=1-\hat{p}_{j}$ remains sufficient in our setting, since we are anyway only concerned with the upper and lower bounds $\bar{V}_{t}$, $\underline{V\mkern-4.0mu}\mkern 4.0mu_{t}$ (the observed mortality being not controllable in practice). Moreover, we also note for the lower bound $\underline{V\mkern-4.0mu}\mkern 4.0mu_{t}$ in Proposition 4.22 (ii) that ${}_{s-t}\underline{p\mkern-4.0mu}\mkern 4.0mu^{*}_{x+t}$ is left-continuous in $s$ and hence not a proper survival function (no minimizer $\underline{\tau\mkern-4.0mu}\mkern 4.0mu^{*}_{j}$). Importantly, the absence of minimizer is again not an issue in our framework, since the observed mortality cannot be controlled in practice. We are in fact only interested in the worst-case contract value $\underline{V\mkern-4.0mu}\mkern 4.0mu_{t}$, which allows the insurer to quantify the maximum potential loss arising from deviations of the realized mortality rates from her chosen mortality assumption/model. If needed, however, ${}_{s-t}\underline{p\mkern-4.0mu}\mkern 4.0mu^{*}_{x+t}$ can always be used in practice as an approximate optimal survival probability.
4.3. Guaranteed Minimum Death Benefits
Guaranteed Minimum Death Benefits (GMDB) ensure that if the policyholder dies before the contract’s maturity, their beneficiaries will receive at least a guaranteed minimum amount, even if the investment’s performance is lower. The payout $\widehat{H}$ in case of death at time $s\geq t$ is then defined similarly as in (3.7) using $G^{D}_{s}=A_{0}e^{r_{g}s}$,
$\widehat{H}(A_{s},G^{D}_{s})=\max\{A_{s},G^{D}_{s}\}=G^{D}_{s}+(A_{s}-G^{D}_{s})_{+}\,$, (4.23)
with $G^{D}_{s}>0$, and whose price $C_{t}(s,G^{D}_{s})$ at time $t\leq s$ is given in Proposition 3.2. The fair value of such a policy can be written using (2.5)–(2.6),
$V^{\mu}_{t}$ $=\mathbb{E}^{\mathbb{Q}}\left[\int_{t}^{T}e^{-\int_{t}^{s}r_{u}du}\,{}_{s^{-}-t}p_{x+t}\left(\mu_{x,c}(s)+\sum_{j\in\mathbb{N}}\mu_{x,j}\,\delta(s-\tau_{j})\right)\widehat{H}(A_{s},G^{D}_{s})\,ds\ \Big|\ \mathcal{H}_{t}\right]$
$\hskip-13.65732pt=\mathbb{E}^{\mathbb{Q}}\hskip-1.99168pt\left[\hskip-1.13809pt\int_{t}^{T}\hskip-3.41432pte^{-\int_{t}^{s}(r_{u}+\mu_{x,c}(u))du}\hskip-1.70717pt\left(\prod_{j\,:\,t<\tau_{j}<s}\hskip-7.11317pt(1-\mu_{x,j})\right)\hskip-2.41849pt\hskip-1.42262pt\left(\hskip-1.42262pt\mu_{x,c}(s)+\sum_{j\in\mathbb{N}}\mu_{x,j}\,\delta(s-\tau_{j})\hskip-2.13394pt\right)\hskip-2.41849pt\widehat{H}(A_{s},G^{D}_{s})\,ds\,\Big|\,\mathcal{H}_{t}\right]$ (4.24)
where $\delta(\cdot)$ is the Dirac measure at zero. We again aim at deriving the upper bound $\bar{V}_{t}=\sup_{\mu\in\mathcal{W}}V^{\mu}_{t}$ and lower bound $\underline{V\mkern-4.0mu}\mkern 4.0mu_{t}=\inf_{\mu\in\mathcal{W}}V^{\mu}_{t}$ subject to the constraint (4.6), where the supremum is taken with respect to both the continuous and the discrete components of the force of mortality.
Proposition 4.5 (GMDB bounds).
Assume that for all $s\in[t,T]$, $r_{g}\geq f(t,s)$ and
$r_{t}\alpha(t)-\theta(t)\big(\alpha(t)+(T-t)\big)\leq 0,\quad\text{with}\quad\alpha(t)=\frac{e^{-\kappa(T-t)}-1}{\kappa}\leq 0.$ (4.25)
(i) Upper bound. The optimal control for $\bar{V}_{t}:=\sup_{\mu\in\mathcal{W}}V_{t}^{\mu}$ is given by
$\bar{\mu}^{*}_{x,c}(t)\equiv 0,\qquad\bar{\mu}^{*}_{x,d}=\{(\bar{\tau}_{j}^{*},\bar{\mu}_{x,j}^{*})\}_{j=0,\ldots,\lfloor T\rfloor-\lceil t\rceil-1},$
with intervention times and jump sizes
$\bar{\tau}_{j}^{*}=\lceil t\rceil+j+1,\qquad\bar{\mu}_{x,j}^{*}=1-\hat{p}_{j}.$
The corresponding upper bound is
$\bar{V}_{t}=\sum_{j=0}^{\lfloor T\rfloor-\lceil t\rceil-1}C_{t}(\lceil t\rceil+j+1,G^{D}_{\lceil t\rceil+j+1})\,(1-\hat{p}_{j})\,\prod_{i<j}\hat{p}_{i}.$ (4.26)
(ii) Lower bound. The lower bound $\underline{V}_{t}:=\inf_{\mu\in\mathcal{W}}V_{t}^{\mu}$ is not attained in $\mathcal{W}$. However, there exists a sequence of admissible controls $\{\mu^{(n)}\}_{n\geq 1}\subset\mathcal{W}$ such that $V_{t}^{\mu^{(n)}}\downarrow\underline{V}_{t}$ and whose discrete intervention times satisfy $\tau^{(n)}_{j}\downarrow\lceil t\rceil+j$ with $\mu^{(n)}_{x,j}=1-\hat{p}_{j}$ for each $j=0,\dots,\lfloor T\rfloor-\lceil t\rceil-1$ and with $\mu_{x,c}^{(n)}\equiv 0$. This lower bound $\underline{V\mkern-4.0mu}\mkern 4.0mu_{t}$ then satisfies
$\underline{V\mkern-4.0mu}\mkern 4.0mu_{t}=\sum_{j=0}^{\lfloor T\rfloor-\lceil t\rceil-1}C_{t}(\lceil t\rceil+j,G^{D}_{\lceil t\rceil+j})\,(1-\hat{p}_{j})\,\prod_{i<j}\,\hat{p}_{i}\,.$ (4.27)
This case differs from the GMIB (4.16), where the policyholder continuously receives benefits while alive. In contrast, we here optimize the single, lump-sump benefit paid at the time of death and thus directly the contract price $C_{t}(\tau,G^{D}_{\tau})$, see the proof A.3. Moreover, Proposition 3.2 provides an analytical formula for $C_{t}(\tau,G^{D}_{\tau})$, which allows us to efficiently compute (4.26) and (4.27). We also note that the assumption $r_{g}\geq f(t,s)$ for all $s\in[t,T]$ is a strong but practically relevant for GMDB contracts and makes this guarantee attractive for risk-averse policyholders. Although this proposition is stated under rather strong assumptions (necessary to obtain analytical expressions), these conditions can be relaxed but the corresponding optimal bounds then have to be computed numerically. Indeed, for more general payoffs $\widehat{H}(A_{s},G^{D}_{s})$ or when the assumptions $r_{g}\geq f(t,s)$ and (4.25) do not hold, one can always use numerical techniques to solve the optimal stopping problems
$\operatorname*{arg\,max/min}_{\tau\in\mathcal{T}_{(\lceil t\rceil+j,\lceil t\rceil+j+1]}}\mathbb{E}^{\mathbb{Q}}\left[e^{-\int_{t}^{\tau}r_{s}ds}\,\widehat{H}(A_{\tau},G^{D}_{\tau})\ \big|\ \mathcal{H}_{t}\right]$ (4.28)
for each $j=0,\ldots,\lfloor T\rfloor-\lceil t\rceil-1$ as it is equivalent to the numerical pricing of $\lfloor T\rfloor-\lfloor t\rfloor$ American options; see Cox et al. (1979) and Longstaff and Schwartz (2001). In the case this optimal intervention time $\tau^{*}_{j}$ solving (4.28) is not unique, we can then again simply choose arbitrarily one such $\tau^{*}_{j}$ as we are only interested in practice in the bounds $\bar{V}_{t}$ (or $\underline{V\mkern-4.0mu}\mkern 4.0mu_{t}$), and not the mortality rate $\mu^{*}$ itself.
As mentioned, the strict matching condition (4.6) with the given life table (2.8) is a strong setting, since in practice the observed mortality over each one-year period $j=0,\ldots,\lfloor T\rfloor-\lceil t\rceil-1$ will deviate from the tabulated $\hat{p}_{j}$. Theoretically, this almost-sure matching of the life tables also restricts the generality of the results in Propositions 4.34.224.5 as we systematically obtain deterministic optimal mortality rates $\mu^{*}$. Section 5 will then relax the set of admissible controls and consider a more general framework, where this strict a.s. matching of the life table will be replaced by a constraint in expectation. This will allow us to consider a more realistic framework where mortality rate trajectories can deviate from the life table while still being consistent on average. Combining Section 4 and 5 will provide robust tools for managing the mortality risk of general variable annuities.
5 Bounds on variable annuities: matching life tables in expectation
In this section, we aim to derive worst- and best-case prices of variable annuities for mortality rate models that remain consistent with the initial life table but under weaker constraints. As mentioned above, we adopt a less restrictive constraint that requires only a matching of the initial life table in expectation. Moreover, the mortality intensity is assumed to be a general progressively measurable process with no explicit discrete component. Therefore, in this section, consistency is ensured by assuming that, for a given age $x$, the mortality intensity $\mu_{x}$ is now a general positive progressively measurable process that satisfies the following constraints from (2.7):
$\mathbb{E}^{\mathbb{P}}\left(e^{-\int_{0}^{j}\mu_{x}(s)ds}\right)={}_{j}\hat{p}_{x},\quad j=1,\cdots,n.$ (5.1)
where $_{j}\hat{p}_{x}$ is again a tabulated value defined by,
$_{j}\hat{p}_{x}:=\prod_{i=0}^{j-1}\hat{p}_{i}$
with $\hat{p}_{i}:={}_{1}p_{x+i}=l_{x+i+1}/l_{x+i}$, see (4.4)–(4.5). In that context, under Assumption 4.2, we aim to solve the following stochastic control problems associated to the worst- and best-case prices defined as [^7] In contrast to the previous section, we adopt a general formulation based on the full combined product, since the dynamic programming approach used to solve the control problem wouldbe essentially the same if each product (GMAB, GMIB, GMDB) were formulated separately. See Section 4 for the definition of each payoff $H$, $\widetilde{H}$, $\widehat{H}$.
$\sup_{\mu_{x}\in\mathcal{A}}\mathbb{E}^{\mathbb{Q}}\Bigg[\,e^{-\int_{0}^{T}(r_{u}+\mu_{x}(u))du}H(A_{T},G^{A}_{T})+\int_{0}^{T}e^{-\int_{0}^{s}(r_{u}+\mu_{x}(u))du}\left(\widetilde{H}(A_{s},G^{I}_{s})+\mu_{x}(s)\widehat{H}(A_{s},G^{D}_{s})\right)ds\Bigg],$ (5.2)
and
$\inf_{\mu_{x}\in\mathcal{A}}\mathbb{E}^{\mathbb{Q}}\Bigg[\,e^{-\int_{0}^{T}(r_{u}+\mu_{x}(u))du}H(A_{T},G^{A}_{T})+\int_{0}^{T}e^{-\int_{0}^{s}(r_{u}+\mu_{x}(u))du}\left(\widetilde{H}(A_{s},G^{I}_{s})+\mu_{x}(s)\widehat{H}(A_{s},G^{D}_{s})\right)ds\Bigg],$ (5.3)
with $T=n$, and
$\small\mathcal{A}:=\left\{\mu_{x}:[0,T]\times\Omega\rightarrow\mathbb{R}^{+}\penalty 10000\ \text{prog. measurable process such that }\mathbb{E}^{\mathbb{P}}\left(e^{-\int_{0}^{j}\mu_{x}(s)ds}\right)={}_{j}\hat{p}_{x},\quad j=1,\cdots,n.\right\}.$ (5.4)
Depending on the structure of the payoff functional, those problems are potentially ill-posed or not tractable. That is why we decide to consider a regularized version of those problems, by restricting the set of admissible controls. For $\delta>0$, we introduce the admissible set of bounded controls $\mathcal{A}^{\delta}_{bounded}$ defined by
$\small\mathcal{A}^{\delta}_{bounded}:=\mathcal{A}\cap\left\{\mu_{x}:[0,T]\times\Omega\rightarrow\mathbb{R}^{+}\penalty 10000\ \text{such that for all }t\leq T,\penalty 10000\ ({\mu_{x,a}}(t)-\delta)_{+}\leq\mu_{x}(t)\leq{\mu_{x,a}}(t)+\delta\penalty 10000\ a.s.\right\},$ (5.5)
with $\mu_{x,a}(t)$ an a priori time-dependent mortality rate consistent with the initial observed life table such that $e^{-\int_{0}^{j}\mu_{x,a}(t)dt}={}_{j}\hat{p}_{x},$ for $j=1,\cdots,n$ [^8] Typically, $\mu_{x,a}$ is obtained by a fractional age assumption.. In this setting, the stochastic control problems become
$\hskip-2.84526pt\sup_{\mu_{x}\in\mathcal{A}^{\delta}_{bounded}}\mathbb{E}^{\mathbb{Q}}\Bigg[\,e^{-\int_{0}^{T}(r_{u}+\mu_{x}(u))du}H(A_{T},G_{T}^{A})+\int_{0}^{T}\hskip-2.84526pte^{-\int_{0}^{s}(r_{u}+\mu_{x}(u))du}\left(\widetilde{H}(A_{s},G^{I}_{s})+\mu_{x}(s)\widehat{H}(A_{s},G_{s}^{D})\right)ds\Bigg],$ (5.6)
and,
$\hskip-2.84526pt\inf_{\mu_{x}\in\mathcal{A}^{\delta}_{bounded}}\mathbb{E}^{\mathbb{Q}}\Bigg[\,e^{-\int_{0}^{T}(r_{u}+\mu_{x}(u))du}H(A_{T},G_{T}^{A})+\int_{0}^{T}\hskip-2.84526pte^{-\int_{0}^{s}(r_{u}+\mu_{x}(u))du}\left(\widetilde{H}(A_{s},G_{s}^{I})+\mu_{x}(s)\widehat{H}(A_{s},G_{s}^{D})\right)ds\Bigg].$ (5.7)
Since we assume that the controls in the set $\mathcal{A}^{\delta}_{bounded}$ are bounded below and above by time-dependent functions consistent with the initial life table then, we can directly deduce that the problems (5.6) and (5.7) are well-posed. Moreover, Proposition 5.4 below shows that as $\delta\to\infty$, the solutions to the regularized problems converge to the corresponding solutions of the original worst-case and best-case pricing problems.
The regularized control problems in (5.6) and (5.7) share similarities with the framework of Li and Szimayer (2011). However, rather than only assuming boundedness of the controlled process, we introduce the expectation constraints (5.1) to ensure consistency with observed life tables. In this sense, the problems are also closely connected to Avellaneda and Paras (1996), who derived bounds on financial derivative prices under volatility uncertainty while maintaining consistency with observed vanilla option prices. For the sake of clarity, we will mainly consider the problem related to the worst-case price, but similar developments can be deduced in the same manner for the problem related to the best case price. Using a dual approach, as done for instance in Motte and Hainaut (2025), we can show that the stochastic control problem (5.6) is equivalent to the following problem (see proof of Proposition 5.2 below):
$\inf_{\boldsymbol{\lambda}\in\mathbb{R}^{n}}\hskip-1.42262pt\sup_{\mu_{x}\in\tilde{\mathcal{A}}^{\delta}_{bounded}}\hskip-5.69054pt\mathbb{E}^{\mathbb{Q}}\Bigg[$ $e^{-\int_{0}^{T}(r_{u}+\mu_{x}(u))du}H(A_{T},G_{T}^{A})\hskip-1.42262pt+\hskip-2.27621pt\int_{0}^{T}\hskip-3.98337pte^{-\int_{0}^{s}(r_{u}+\mu_{x}(u))du}\hskip-1.42262pt\left(\widetilde{H}(A_{s},G_{s}^{I})\hskip-1.42262pt+\hskip-1.42262pt\mu_{x}(s)\widehat{H}(A_{s},G_{s}^{D})\hskip-1.42262pt\right)ds\Bigg]$ (5.8)
$-\sum_{j=1}^{n}\lambda_{j}\left[\mathbb{E}^{\mathbb{P}}\left(e^{-\int_{0}^{j}\mu_{x}(s)ds}\right)-{}_{j}\hat{p}_{x}\right],$
with $\boldsymbol{\lambda}:=(\lambda_{1},\cdots,\lambda_{n})\in\mathbb{R}^{n}$ and, for $\delta>0$,
$\tilde{\mathcal{A}}^{\delta}_{bounded}:=\left\{\mu_{x}:[0,T]\times\Omega\rightarrow\mathbb{R}^{+}\penalty 10000\ \text{s.t.\ for all }t\leq T,\penalty 10000\ ({\mu_{x,a}}(t)-\delta)_{+}\leq\mu_{x}(t)\leq{\mu_{x,a}}(t)+\delta\penalty 10000\ a.s.\right\}.$
Let us show how to solve the inner problem using a dynamic programming approach. To this end, we assume that the unique change of measure process $(Z_{t})_{0\leq t\leq T}$ defined by (3.2) is given by
$Z_{t}:=\exp\left(-\int_{0}^{t}\left(\frac{\mu_{S}-r_{s}}{\sigma_{S}}\right)d{W}_{s}^{1,\mathbb{P}}-\frac{1}{2}\int_{0}^{t}\left(\frac{\mu_{S}-r_{s}}{\sigma_{S}}\right)^{2}ds\right),\quad 0\leq t\leq T,$ (5.9)
with $\mu_{S}$ the drift of the stock under the real measure $\mathbb{P}$ and ${W}^{1,\mathbb{P}}$ a standard Brownian motion defined on $\mathbb{P}$ such that
${W}^{1,\mathbb{P}}_{t}:=W^{1,\mathbb{Q}}_{t}-\int_{0}^{t}\left(\frac{\mu_{S}-r_{s}}{\sigma_{S}}\right)ds,\quad 0\leq t\leq T,$
with ${W}^{1,\mathbb{Q}}$ the standard Brownian motion defined on $\mathbb{Q}$ from Section 3.
Remark 5.1.
In our framework, we consider a complete financial market, which guarantees the uniqueness of the risk-neutral measure $\mathbb{Q}$. Under $\mathbb{Q}$, the stock price $(S_{t})_{0\leq t\leq T}$ has a drift equal to the short rate $(r_{t})_{0\leq t\leq T}$, as required for arbitrage-free pricing. For the short rate $(r_{t})_{0\leq t\leq T}$, the change of measure from the physical measure $\mathbb{P}$ to $\mathbb{Q}$ generally involves a market price of risk $(\lambda^{r}_{t})_{0\leq t\leq T}$, which is calibrated in practice to match observed prices of traded instruments, such as zero-coupon bonds. For simplicity and consistency with Section 4, we assume that $\lambda^{r}_{t}=0$ for all $0\leq t\leq T$ but other choices of $(\lambda^{r}_{t})_{0\leq t\leq T}$ could be made.
Let us define the process $\tilde{Z}=(\tilde{Z}_{t})_{0\leq t\leq T}$ as
$\tilde{Z}_{t}:=e^{-\int_{0}^{t}r_{s}ds}Z_{t},\quad 0\leq t\leq T,$ (5.10)
with $Z$ defined by (5.9). For $\boldsymbol{\lambda}\in\mathbb{R}^{n}$, the inner problem in (5.8) can be rewritten as
$\sup_{\mu_{x}\in\tilde{\mathcal{A}}^{\delta}_{bounded}}\mathbb{E}^{\mathbb{Q}}\Bigg[$ $e^{-\int_{0}^{T}(r_{u}+\mu_{x}(u))du}H(A_{T},G_{T}^{A})-\sum_{j=1}^{n}\frac{\lambda_{j}}{\tilde{Z}_{j}}e^{-\int_{0}^{j}(r_{s}+\mu_{x}(s))ds}$ (5.11)
$+\int_{0}^{T}e^{-\int_{0}^{s}(r_{u}+\mu_{x}(u))du}\left(\widetilde{H}(A_{s},G_{s}^{I})+\mu_{x}(s)\widehat{H}(A_{s},G_{s}^{D})\right)ds\Bigg].$
To solve this control problem, for $\boldsymbol{\lambda}\in\mathbb{R}^{n}$, we introduce the following value function
$\bar{v}(t,a,r,z;\boldsymbol{\lambda})=\hskip-2.84526pt\sup_{\mu_{x}\in\tilde{\mathcal{A}}^{\delta}_{bounded}(t,T)}\hskip-2.84526pt$ $\mathbb{E}_{t,a,r,z}^{\mathbb{Q}}\Bigg[\,e^{-\int_{t}^{T}(r_{u}+\mu_{x}(u))du}F(A_{T},G_{T}^{A})-\sum_{j=\lceil t\rceil}^{n}\frac{\lambda_{j}}{\tilde{Z}_{j}}e^{-\int_{t}^{j}(r_{s}+\mu_{x}(s))ds}$
$\hskip-0.56905pt+\hskip-1.42262pt\int_{t}^{T}\hskip-3.41432pte^{-\int_{t}^{s}(r_{u}+\mu_{x}(u))du}\hskip-1.42262pt\left(\widetilde{H}(A_{s},G_{s}^{I})+\mu_{x}(s)H(A_{s},G_{s}^{D})\right)\hskip-0.85358ptds\Bigg],\hskip-2.84526pt$ (5.12)
where $\mathbb{E}^{\mathbb{Q}}_{t,a,r,z}(.)$ denotes the conditional expectation given $A_{t}=a$, $r_{t}=r$, $\tilde{Z}_{t}=z$, and
$\small\mathcal{A}^{\delta}_{bounded}(t,T):=\left\{\mu_{x}:[t,T]\times\Omega\rightarrow\mathbb{R}^{+}\penalty 10000\ \text{s.t.\ for all }s\in[t,T],\penalty 10000\ ({\mu_{x,a}}(s)-\delta)_{+}\leq\mu_{x}(t)\leq{\mu_{x,a}}(s)+\delta\penalty 10000\ a.s.\right\}.$ (5.13)
Let us define $A^{\delta}(t)$ as
$A^{\delta}(t):=[(\mu_{x,a}(t)-\delta)_{+},\mu_{x,a}(t)+\delta],\quad t\leq T.$ (5.14)
Classical results from dynamic programming (see for instance Touzi (2012)) suggest that the value function associated to the inner problem solves the following Hamilton-Jacobi-Bellman (HJB) equation, for $t\in({j-1},j)$, j=1,,nj=1,...,n:
$\sup_{\mu_{x}\in A^{\delta}(t)}\left[\mathcal{L}_{t,a,r,z}\bar{v}(t,a,r,z;\boldsymbol{\lambda})+\widetilde{H}(a,G_{t}^{I})+\mu_{x}\left(\widehat{H}(a,G_{t}^{D})-\bar{v}(t,a,r,z;\boldsymbol{\lambda})\right)\right]=0,$ (5.15)
with boundary conditions
$\bar{v}(n,a,r,z;\boldsymbol{\lambda})=H(a,G_{t_{n}}^{A})-\frac{\lambda_{n}}{z},$ (5.16)
$\bar{v}({j}^{-},a,r,z;\boldsymbol{\lambda})=\bar{v}({j}^{+},a,r,z;\boldsymbol{\lambda})-\frac{\lambda_{j}}{z},\quad j=1,.,n-1,$
where $\mathcal{L}_{t,a,r,z}$ is the classical generator associated with $(A_{t})_{0\leq t\leq T}$ given by (3.6), $(r_{t})_{0\leq t\leq T}$ given by (3.1), and $(\tilde{Z}_{t})_{0\leq t\leq T}$ given by (5.10). Since (5.15) depends linearly on $\mu$, we deduce that, for $j=1,\cdots,n$ and $t\in({j-1},{j})$,
$\sup_{\mu_{x}\in A^{\delta}(t)}$ $\mu_{x}\left(\widehat{H}(a,G_{t}^{D})-\bar{v}(t,a,r,z;\boldsymbol{\lambda})\right)=$
[…]
Therefore, the maximizer $\bar{\mu}_{x}^{*}$ is given by, for $j=1,\cdots,n$ and $t\in({j-1},{j})$,
${\bar{\mu}}_{x}^{*}(t,a,r,z;\boldsymbol{\lambda})=\begin{cases}(\mu_{x,a}(t)-\delta)_{+},&\text{if }\widehat{H}(a,G_{t}^{D})<\bar{v}(t,a,r,z;\boldsymbol{\lambda}),\\ (\mu_{x,a}(t)+\delta),&\text{if }\widehat{H}(a,G_{t}^{D})\geq\bar{v}(t,a,r,z;\boldsymbol{\lambda}).\end{cases}$ (5.17)
Plugging this into the HJB equation (5.15), we deduce that candidate value function satisfies the following nonlinear PDE:
$\mathcal{L}_{t,a,r,z}\bar{v}(t,a,r,z;\boldsymbol{\lambda})+\widetilde{H}(a,G_{t}^{I})+{\bar{\mu}}_{x}^{*}(t,a,r,z;\boldsymbol{\lambda})\left(\widehat{H}(a,G_{t}^{D})-\bar{v}(t,a,r,z;\boldsymbol{\lambda})\right)=0,$ (5.18)
with boundary conditions given by (5.16).
Proposition 5.2 shows that under mild regularity assumptions, the candidate value function solution of the HJB equation is effectively the solution to the regularized inner control problem associated to the worst case price. Proposition 5.3 derives an analogue result for the the best case problem. Finally, Proposition 5.4 shows that the solutions of the regularized problems converge toward the solutions of the original worst- and best-case problems given by (5.2) and (5.3). Proofs are once again deferred in A.
Proposition 5.2.
For any $\bm{\lambda}\in\mathbb{R}^{n}$, let assume that there exists $\bar{v}(t,a,r,z;\boldsymbol{\lambda})\in C^{1,2,2,2}$ solution to (5.18). Then, $\bar{v}(t,a,r,z;\boldsymbol{\lambda})$ solves the stochastic control problem (5), and the associated optimal mortality rate is given by (5.17). In particular, for $\boldsymbol{\lambda}\in\mathbb{R}^{n}$, $\bar{v}(0,A_{0},r_{0},Z_{0};\boldsymbol{\lambda})$ solves the initial inner problem such that
$\bar{v}(0,A_{0},r_{0},Z_{0};\boldsymbol{\lambda})$
$=\hskip-1.42262pt\sup_{\mu_{x}\in\tilde{\mathcal{A}}^{\delta}_{bounded}}\hskip-2.84526pt\mathbb{E}^{\mathbb{Q}}\Bigg[e^{-\int_{0}^{T}(r_{u}+\mu_{x}(u))du}H(A_{T},G_{T}^{A})+\hskip-1.99168pt\int_{0}^{T}\hskip-2.84526pte^{-\int_{0}^{s}(r_{u}+\mu_{x}(u))du}\hskip-1.13809pt\left(\widetilde{H}(A_{s},G_{s}^{I})+\mu_{x}(s)\widehat{H}(A_{s},G_{s}^{D})\right)\hskip-1.13809ptds\Bigg]$
$\hskip 56.9055pt-\sum_{j=1}^{n}\lambda_{j}\mathbb{E}^{\mathbb{P}}\left(e^{-\int_{0}^{j}\mu_{x}(s)ds}\right).$
Moreover, the infimum of $\boldsymbol{\lambda}\mapsto\left(\bar{v}(0,A_{0},r_{0},Z_{0};\boldsymbol{\lambda})+\sum_{j=1}^{n}\lambda_{j}\,{}_{j}\hat{p}_{x}\right)$ over $\boldsymbol{\lambda}\in\mathbb{R}^{n}$ is attained at some $\boldsymbol{\lambda}^{*}\in\mathbb{R}^{n}$; hence
$\bar{v}_{*}(0,A_{0},r_{0},Z_{0}):=\bar{v}(0,A_{0},r_{0},Z_{0};\boldsymbol{\lambda}^{*})+\sum_{j=1}^{n}\lambda^{*}_{j}\,{}_{j}\hat{p}_{x}=\min_{\boldsymbol{\lambda}\in\mathbb{R}^{n}}\left[\bar{v}(0,A_{0},r_{0},Z_{0};\boldsymbol{\lambda})+\sum_{j=1}^{n}\lambda_{j}\,{}_{j}\hat{p}_{x}\right],$ (5.19)
is well-defined and is the solution of the regularized worst-case price problem
$\bar{v}_{*}(0,A_{0},r_{0},Z_{0})=\sup_{\mu_{x}\in\mathcal{A}^{\delta}_{bounded}}\mathbb{E}^{\mathbb{Q}}\Bigg[$ $\,e^{-\int_{0}^{T}(r_{u}+\mu_{x}(u))du}H(A_{T},G_{T}^{A})$ (5.20)
$+\int_{0}^{T}e^{-\int_{0}^{s}(r_{u}+\mu_{x}(u))du}\left(\widetilde{H}(A_{s},G_{s}^{I})+\mu_{x}(s)\widehat{H}(A_{s},G_{s}^{D})\right)ds\Bigg].$
Proposition 5.3.
For any $\bm{\lambda}\in\mathbb{R}^{n}$, let assume that, there exists $\underline{v\mkern-4.0mu}\mkern 4.0mu(t,a,r,z;\boldsymbol{\lambda})\in C^{1,2,2,2}$ solution to the following nonlinear PDE, for $t\in({j-1},j),$ j=1,,nj=1,...,n,
$\mathcal{L}_{t,a,r,z}\underline{v\mkern-4.0mu}\mkern 4.0mu(t,a,r,z;\boldsymbol{\lambda})+\widetilde{H}(a,G_{t}^{I})+{\underline{\mu\mkern-4.0mu}\mkern 4.0mu}_{x}^{*}(t,a,r,z;\boldsymbol{\lambda})\left(\widehat{H}(a,G_{t}^{D})-\underline{v\mkern-4.0mu}\mkern 4.0mu(t,a,r,z;\boldsymbol{\lambda})\right)=0,$ (5.21)
with
${\underline{\mu\mkern-4.0mu}\mkern 4.0mu}_{x}^{*}(t,a,r,z;\boldsymbol{\lambda})=\begin{cases}(\mu_{x,a}(t)-\delta)_{+},&\text{if }\widehat{H}(a,G_{t}^{D})>\underline{v\mkern-4.0mu}\mkern 4.0mu(t,a,r,z;\boldsymbol{\lambda}),\\ (\mu_{x,a}(t)+\delta),&\text{if }\widehat{H}(a,G_{t}^{D})\leq\underline{v\mkern-4.0mu}\mkern 4.0mu(t,a,r,z;\boldsymbol{\lambda}),\end{cases}$ (5.22)
that satisfies the following boundary conditions
$\underline{v\mkern-4.0mu}\mkern 4.0mu(n,a,r,z;\boldsymbol{\lambda})=H(a,G_{t_{n}}^{A})+\frac{\lambda_{n}}{z},$ (5.23)
$\underline{v\mkern-4.0mu}\mkern 4.0mu(j^{-},a,r,z;\boldsymbol{\lambda})=\underline{v\mkern-4.0mu}\mkern 4.0mu(j^{+},a,r,z;\boldsymbol{\lambda})+\frac{\lambda_{j}}{z},\quad j=1,.,n-1.$
Then, for $\boldsymbol{\lambda}\in\mathbb{R}^{n}$, $\underline{v\mkern-4.0mu}\mkern 4.0mu(0,A_{0},r_{0},Z_{0};\boldsymbol{\lambda})$ solves the initial inner problem such that
$\underline{v\mkern-4.0mu}\mkern 4.0mu(0,A_{0},r_{0},Z_{0};\boldsymbol{\lambda})$
$=\hskip-1.99168pt\inf_{\mu_{x}\in\tilde{\mathcal{A}}^{\delta}_{bounded}}\hskip-3.69885pt\mathbb{E}^{\mathbb{Q}}\Bigg[e^{-\int_{0}^{T}(r_{u}+\mu_{x}(u))du}H(A_{T},G_{T}^{A})+\hskip-1.42262pt\int_{0}^{T}\hskip-3.1298pte^{-\int_{0}^{s}(r_{u}+\mu_{x}(u))du}\hskip-0.7113pt\left(\widetilde{H}(A_{s},G_{s}^{I})+\mu_{x}(s)\widehat{H}(A_{s},G_{s}^{D})\right)\hskip-0.7113ptds\Bigg]$
$\hskip 56.9055pt+\sum_{j=1}^{n}\lambda_{j}\mathbb{E}^{\mathbb{P}}\left(e^{-\int_{0}^{j}\mu_{x}(s)ds}\right).$
Moreover, the supremum of $\boldsymbol{\lambda}\mapsto\left(\underline{v\mkern-4.0mu}\mkern 4.0mu(0,A_{0},r_{0},Z_{0};\boldsymbol{\lambda})-\sum_{j=1}^{n}\lambda_{j}\,{}_{j}\hat{p}_{x}\right)$ over $\boldsymbol{\lambda}\in\mathbb{R}^{n}$ is attained at some $\boldsymbol{\lambda}^{*}\in\mathbb{R}^{n}$; hence
$\underline{v\mkern-4.0mu}\mkern 4.0mu_{*}(0,A_{0},r_{0},Z_{0}):=\underline{v\mkern-4.0mu}\mkern 4.0mu(0,A_{0},r_{0},Z_{0};\boldsymbol{\lambda}^{*})-\sum_{j=1}^{n}\lambda^{*}_{j}\,{}_{j}\hat{p}_{x}=\max_{\boldsymbol{\lambda}>0}\left[\underline{v\mkern-4.0mu}\mkern 4.0mu(0,A_{0},r_{0},Z_{0};\boldsymbol{\lambda})-\sum_{j=1}^{n}\lambda_{j}{}_{j}\hat{p}_{x}\right],$ (5.24)
is well-defined and is the solution of the regularized best-case price problem
$\underline{v\mkern-4.0mu}\mkern 4.0mu_{*}(0,A_{0},r_{0},Z_{0})=\inf_{\mu_{x}\in\mathcal{A}^{\delta}_{bounded}}\mathbb{E}^{\mathbb{Q}}\Bigg[$ $\,e^{-\int_{0}^{T}(r_{u}+\mu_{x}(u))du}H(A_{T},G_{T}^{A})$ (5.25)
$+\int_{0}^{T}e^{-\int_{0}^{s}(r_{u}+\mu_{x}(u))du}\left(\widetilde{H}(A_{s},G_{s}^{I})+\mu_{x}(s)\widehat{H}(A_{s},G_{s}^{D})\right)ds\Bigg].$
Proof.
The proof follows by arguments analogous to the proof of Proposition 5.2. ∎
Proposition 5.4.
Let us assume that the assumptions of Proposition 5.2 and 5.3 are satisfied. For $\delta>0$, let $\bar{v}^{\delta}_{*}(0,A_{0},r_{0},Z_{0})$ and $\underline{v\mkern-4.0mu}\mkern 4.0mu^{\delta}_{*}(0,A_{0},r_{0},Z_{0})$ denote the solution to the worst- and best-case price problems given by (5.6) and (5.7). Moreover, let $\tilde{\mathcal{A}}$ be the set of bounded controls belonging to $\mathcal{A}$. Then, we have that
$\sup_{\mu_{x}\in\tilde{\mathcal{A}}}\mathbb{E}^{\mathbb{Q}}\Bigg[e^{-\int_{0}^{T}(r_{u}+\mu_{x}(u))du}H(A_{T},G_{T}^{A})+\hskip-1.42262pt\int_{0}^{T}\hskip-2.84526pte^{-\int_{0}^{s}(r_{u}+\mu_{x}(u))du}\hskip-1.13809pt\left(\widetilde{H}(A_{s},G_{s}^{I})+\mu_{x}(s)\widehat{H}(A_{s},G_{s}^{D})\right)\hskip-1.13809ptds\Bigg]$ (5.26)
$\hskip 14.22636pt=\lim_{\delta\to\infty}\bar{v}^{\delta}_{*}(0,A_{0},r_{0},Z_{0}),$
and
$\inf_{\mu_{x}\in\tilde{\mathcal{A}}}\mathbb{E}^{\mathbb{Q}}\Bigg[e^{-\int_{0}^{T}(r_{u}+\mu_{x}(u))du}H(A_{T},G_{T}^{A})+\hskip-1.42262pt\int_{0}^{T}\hskip-3.41432pte^{-\int_{0}^{s}(r_{u}+\mu_{x}(u))du}\hskip-1.13809pt\left(\widetilde{H}(A_{s},G_{s}^{I})+\mu_{x}(s)\widehat{H}(A_{s},G_{s}^{D})\right)\hskip-1.13809ptds\Bigg]$ (5.27)
$\hskip 14.22636pt=\lim_{\delta\to\infty}\underline{v\mkern-4.0mu}\mkern 4.0mu^{\delta}_{*}(0,A_{0},r_{0},Z_{0}).$
6 Numerical experiments
Since our method in Section 4 and 5 does not rely on any specific mortality modeling assumptions, the resulting bounds provided below are the most general one can obtain using the available information (provided that the mortality rates over each one-year period are consistent with the given table). These bounds are therefore of high practical interest for insurers, as they naturally provide a range within which the contract price should lie. Any price outside this range is either not consistent with the life tables or with the specified financial market. Finally, these results allow the insurer to directly quantify its worst-case loss or best-case gain in case of deviation of the observed mortality from the chosen fractional-age/mortality rate assumption. The results are based on the Belgian female life table.
6.1. Strict bounds
The parameters used in this section are $A_{0}=100$, $r_{g}=0.03$, $r_{0}=0.01$, $\pi_{S}=0.6$, $\pi_{P}=0.2$, $\sigma_{S}=0.2$, $\sigma_{r}=0.02$, $\rho=0.3$, and $\kappa=0.1$. The mean-reversion interest rate level $\theta(t)$ is obtained from (3.3) by specifying the instantaneous forward rate as
$f(0,t)=b+ce^{-at}+date^{-at},\qquad a,b,c,d\in\mathbb{R},$
with yield-curve parameters $a=0.015$, b=0.0105b=-0.0105, $c=0.02$, and $d=0.75$. We first depict in Figure 6.1 the survival probabilities $_{t}p_{x}$ at age $x=40$ and $x=100$ for $t\in[0,5]$. Naturally, at lower age $x$, we see that the specific choice of fractional age assumption has almost no impact on the survival probabilities, while we observe higher discrepancies for $x=100$. By contrast, the GMIB- and GMDB-induced optimal survival curves $_{t}\bar{p}^{*}_{x}$ and $_{t}\underline{p\mkern-4.0mu}\mkern 4.0mu^{*}_{x}$, given by (4.20) and (4.22), deviate significantly from all the fractional age assumptions.
Figure 6.1: Comparison of survival probabilities $_{t}p_{x}$ at age $x=40$ (left) and $x=100$ (right) over time $t\in[0,5]$, using the fractional age assumptions (2.9)-(2.11) and optimal GMIB (and GMDB) $_{t}\bar{p}^{*}_{x}$ (’Sup’), $_{t}\underline{p\mkern-4.0mu}\mkern 4.0mu^{*}_{x}$ (’Inf’) from Eqs. (4.20)–(4.22).
We then study the strict bounds of the GMIB contract in Figure 6.2.
Figure 6.2: GMIB contract price $V^{\mu}_{t}$ at $t=0$ for different maturities $T$ with initial age $x=40$ (top left), $x=60$ (top right) and $x=80$ (bottom). Results are shown under the fractional-age assumptions (2.9)–(2.11), together with the GMIB upper bound $\bar{V}_{0}$ (’Sup’) and lower bound $\underline{V\mkern-4.0mu}\mkern 4.0mu_{0}$ (’Inf’) from Eqs. (4.19)–(4.21).
For younger policyholders (smaller initial age $x$), the upper-left plot shows that the insurer’s exposure to deviations of mortality within each one-year period from the chosen fractional-age assumption is moderate, at least for relatively short maturities $T$ (while remaining consistent with the tabulated survival probabilities $\hat{p}_{j}$’s). In contrast, for older policyholders (bottom plot) or for longer maturities, this mortality risk becomes much more significant, as deviations within each one-year period from the fractional-age assumption lead to a stronger impact on the GMIB value. We also observe that the choice among the three fractional-age assumptions (2.9)–(2.11) has virtually no impact on the contract price. All these observations are confirmed for the GMDB bounds in Figure 6.3. Bounds for the GMAB contract are not depicted as they are trivial in the strict pathwise setting, see Proposition 4.3.
Figure 6.3: GMDB contract price $V^{\mu}_{t}$ at $t=0$ for different maturities $T$ with initial age $x=40$ (top left), $x=60$ (top right) and $x=80$ (bottom). Results are shown under the fractional-age assumptions (2.9)–(2.11), together with the GMDB upper bound $\bar{V}_{0}$ (’Sup’) and lower bound $\underline{V\mkern-4.0mu}\mkern 4.0mu_{0}$ (’Inf’) from Eqs. (4.26)–(4.27).
6.2. Relaxed bounds
For clarity, in this section we consider a deterministic interest rate $r=0.01$ and set the market price of risk to zero ($\mu_{S}=r$), which reduces the Hamilton–Jacobi–Bellman (HJB) equations from four state variables to only two. While this simplification ignores stochastic interest rate effects, the qualitative behavior of bounds is expected to remain similar in the full four-state formulation. The other parameters used in the numerical examples are the same as in Section 6.1 above with $\delta=1$. Moreover, the a priori time-dependent mortality rate $\mu_{x,a}$ is given by the Balducci approximation. We compare different variable annuities prices:
  • Prices obtained by using the Balducci approximation,
  • Worst- and best-case prices by imposing expected constraints only at maturity $T$,
  • Worst- and best-case prices by imposing expected constraints for each intermediate maturities $j\leq T$,
  • Worst- and best-case prices by imposing almost surely constraints for each intermediate maturities $j\leq T$.
Figures 6.4, 6.5 and 6.6 display the upper and lower bounds (worst- and best-cases prices) for GMAB, GMIB, and GMDB for policyholder ages $x=40,60,80$. Comparing expected bounds with almost-sure bounds, it is clear that the expected bounds are less restrictive, as they only enforce average behavior over time rather than across all possible mortality trajectories. Consequently, the expected bounds are always wider than the almost-sure bounds. Note that for GMAB, the almost-sure bounds coincide with the price under Balducci approximation, so they are not displayed; for GMIB and GMDB, the almost-sure bounds are tighter and lie closer to the baseline.
Moreover, we observe that adding additional constraints at intermediate times further narrows the bounds, as they restrict extreme mortality trajectories that would otherwise satisfy the final maturity condition but produce very high or very low liabilities along the path. The effect of these intermediate-time constraints is stronger for GMIB and GMDB than for GMAB because GMIB and GMDB payoffs depend on survival or death at intermediate times, whereas GMAB depends primarily on survival at maturity $T$.
We also observe asymmetry in the bounds, when comparing upper and lower bounds to the baseline. This asymmetry depends on the nature of the payoff. For survival-contingent products such as GMIB and GMAB, the lower bounds are typically far below the baseline, reflecting the strong impact of higher mortality on survival-dependent liabilities, while the upper bounds remain closer. For death-contingent products such as GMDB, the opposite pattern is observed: the upper bounds lie far above the baseline, as higher mortality increases the present value of death benefits, whereas the lower bounds remain near the baseline, since reductions in mortality have limited effect. Those asymmetrical effects are reduced for older ages, as survival probabilities decrease.
Figure 6.4: GMAB contract price for different maturities $T\in\{1,2,3,5,7,10,15\}$ with initial age $x=40$ (top left), $x=60$. (top right) and $x=80$ (bottom). Results are shown under the Balducci assumption (2.11), together with the optimal relaxed bounds from Proposition 5.25.3
Figure 6.5: GMIB contract price for different maturities $T\in\{1,2,3,5,7,10,15\}$ with initial age $x=40$ (top left), $x=60$. (top right) and $x=80$ (bottom). Results are shown under the Balducci assumption (2.11), together with the optimal relaxed bounds from Proposition 5.25.3 and a.s. bounds from Eqs. (4.19)–(4.21)
Figure 6.6: GMDB contract price for different maturities $T\in\{1,2,3,5,7,10,15\}$ with initial age $x=40$ (top left), $x=60$. (top right) and $x=80$ (bottom). Results are shown under the Balducci assumption (2.11), together with the optimal relaxed bounds from Proposition 5.25.3 and a.s. bounds from Eqs. (4.26)–(4.27).
We also check the convergence of the upper and lower bounds with expected constraints only at maturity, as $\delta\to\infty$. Figure 6.7 displays the upper and lower bond prices for $x=60$ and $T=5$, plotted as functions of $\delta>0$. As expected, we observe a relative fast convergence as $\delta$ increases, confirming the theoretical result of Proposition 5.4.
Figure 6.7: Optimal relaxed bounds from Proposition 5.25.3 with expected constraint only at maturity $T$ as functions of $\delta>0$, for GMAB, GMIB and GMDB contract prices with initial age $x=60$ and maturity $T=5$.
We conclude this numerical section with some remarks on the computational aspects of the problem and the methodological choices adopted in its resolution. The original control problems are formulated as a four-dimensional HJB equations with multiple time constraints on $[0,T]$. This leads to a sequence of recursive HJB equations, which becomes computationally very demanding, especially for large maturities. In addition, the numerical resolution requires solving a high-dimensional static optimization problem with respect to $\boldsymbol{\lambda}\in\mathbb{R}^{n}$, which further increases the computational burden and slows convergence. To ensure tractable computation and reliable numerical results, we decide to reduce the state space to two variables. The full four-dimensional formulation, with stochastic interest rates, could in principle be addressed using modern deep-learning-based methods for high-dimensional control problems instead of the finite-difference scheme used above, see for example Cheridito et al. (2025), Dupret and Hainaut (2026), which are specifically designed to handle such large systems of recursive HJB equations. Let us also note that the previous section provides alternative almost-sure bounds obtained under stronger and less realistic assumptions. Although more restrictive, these bounds are fast to compute and can be derived directly. They thus serve as useful a priori estimates and sanity checks for the numerical results.
7 Conclusion
This paper addresses a structural source of model risk in life insurance within a joint actuarial-financial framework: life tables provide survival probabilities only at integer ages, leaving the within-year timing of deaths unspecified, even though many modern life-contingent liabilities (variable annuities) depend on the full continuous-time lifetime distribution. Rather than postulating a fractional-age assumption (UDD/CFM/Balducci) or a parametric (deterministic or stochastic) mortality rate model, we develop a model-free mortality framework that characterizes the set of admissible mortality intensities compatible with the tabulated one-year survival probabilities and the given financial market, yielding sharp upper and lower bounds for hybrid actuarial-financial functionals over that set. We present two distinct yet complementary approaches. In the strict (pathwise) matching regime, admissible mortality trajectories must reproduce each observed one-year survival probability almost surely. Because strict matching is often too restrictive empirically and lead to deterministic optimal mortality rates, we then study a relaxed regime where life-table constraints are matched in expectation. To avoid the control problem to be ill-posed, we introduce a regularized admissible set around a baseline mortality intensity consistent with the table, and solve the resulting robust control problem via a dual formulation with Lagrange multipliers and dynamic programming. Numerical experiments confirm intuitive patterns: (i) expectation-based bounds are wider than almost-sure bounds; (ii) adding intermediate-time constraints tightens bounds substantially for GMIB/GMDB; and (iii) the bound asymmetry depends on whether the payoff is survival- or death-contingent, with the effect diminishing at very old ages as survival probabilities shrink. Overall, the proposed bounds provide a practical, model-free mortality risk-management layer on top of any chosen financial model.
Appendix A Proofs
A.1. Proof of Proposition 4.3
Proof.
Using (4.9), (4.6) and (4.5), we have that for every $\mu\in\mathcal{W}$,
$V^{\mu}_{t}$ $=\mathbb{E}^{\mathbb{Q}}\left[\left(\hskip-0.28453pt\prod_{j=0}^{\lfloor T\rfloor-\lceil t\rceil-1}\hskip-2.84526pt\hat{p}_{j}\right)e^{-\int_{t}^{T}r_{s}ds}\,H(A_{T},G^{A}_{T})\ \big|\ \mathcal{H}_{t}\right]$
$={}_{T-t}p_{x+t}\,\mathbb{E}^{\mathbb{Q}}\bigg[\,e^{-\int_{t}^{T}r_{s}ds}\,H(A_{T},G^{A}_{T})\ \big|\ \mathcal{H}_{t}\,\bigg]\,.$
Under Assumption 4.2, $\mathbb{E}^{\mathbb{Q}}[X\,|\,\mathcal{H}_{t}]=\mathbb{E}^{\mathbb{Q}}[X\,|\,\mathcal{F}_{t}]$ for any $\mathcal{F}_{T}$-measurable integrable $X$. Then,
$V^{\mu}_{t}={}_{T-t}p_{x+t}\,C_{t}(T,G^{A}_{T})\,.$
As the survival probability $_{T-t}p_{x+t}$ is known and given by (4.5), the objective functional $V^{\mu}_{t}$ does not depend on the underlying mortality control $\mu\in\mathcal{W}$. Therefore, any admissible $\mu$ is a maximizer and minimizer of $V^{\mu}_{t}$ and we find $\bar{V}_{t}=\underline{V\mkern-4.0mu}\mkern 4.0mu_{t}=V^{\mu}_{t}$. ∎
A.2. Proof of Proposition 4.22
Proof.
(i) Upper bound. Fix the interval $(\lceil t\rceil+j,\lceil t\rceil+j+1]$ for some $j=0,\ldots,\lfloor T\rfloor-\lceil t\rceil-1$. The life-table constraint (4.6) forces the total survival factor over $(\lceil t\rceil+j,\lceil t\rceil+j+1]$ to be $\hat{p}_{j}$. Since $\widetilde{H}\geq 0$, the integrand
$s\mapsto e^{-\int_{t}^{s}r_{u}\,du}\,\widetilde{H}(A_{s},G_{s}^{I})$
is nonnegative. Let $\mathcal{T}_{(a,b]}$ denote the set of $\mathbb{H}$-stopping times on $(a,b]$. Therefore, for any $\tau_{1},\tau_{2}\in\mathcal{T}_{(\lceil t\rceil+j,\lceil t\rceil+j+1]}$ with $\tau_{1}\leq\tau_{2}$ a.s.,
$\int_{a}^{\tau_{1}}e^{-\int_{t}^{s}r_{u}\,du}\,\widetilde{H}(A_{s},G_{s}^{I})\,ds\leq\int_{a}^{\tau_{2}}e^{-\int_{t}^{s}r_{u}\,du}\,\widetilde{H}(A_{s},G_{s}^{I})\,ds\,,\quad\text{a.s.}$
Taking conditional expectations given $\mathcal{H}_{t}$ leads to
$\mathbb{E}^{\mathbb{Q}}\left[\int_{\lceil t\rceil+j}^{\tau_{1}}e^{-\int_{t}^{s}r_{u}du}\,\widetilde{H}(A_{s},G^{I}_{s})\,ds\ \big|\ \mathcal{H}_{t}\right]\leq\mathbb{E}^{\mathbb{Q}}\left[\int_{\lceil t\rceil+j}^{\tau_{2}}e^{-\int_{t}^{s}r_{u}du}\,\widetilde{H}(A_{s},G^{I}_{s})\,ds\ \big|\ \mathcal{H}_{t}\right].$ (A.1)
Hence the conditional expected accumulated discounted payoff (A.1) over $(\lceil t\rceil+j,\tau]$ is maximized by taking $\tau=\lceil t\rceil+j+1$. Consequently, we have from (4.16) that $V_{t}^{\mu}$ is maximized when all the mortality mass available on the interval $(\lceil t\rceil+j,\lceil t\rceil+j+1]$ is assigned at time $\lceil t\rceil+j+1$. This is achieved by setting $\bar{\mu}_{x,c}^{*}\equiv 0$ (since continuous process with values in a compact set) and choosing a single discrete intervention on each $(\lceil t\rceil+j,\lceil t\rceil+j+1]$ at $\bar{\tau}_{j}^{*}=\lceil t\rceil+j+1$. The life-table constraint (4.6) then directly leads to the optimal jump size $\bar{\mu}_{x,j}^{*}=1-\hat{p}_{j}$.
We can hence rewrite the optimal survival probability
${}_{s-t}\bar{p}^{*}_{x+t}=\prod_{j\,:\,\bar{\tau}^{*}_{j}\leq s}\hat{p}_{j}=\prod_{j\,:\,\lceil t\rceil+j+1\leq s}\hat{p}_{j}\,$,
with $\prod_{\emptyset}:=1$. As $\bar{\mu}^{*}$ (and hence $_{s-t}\bar{p}^{*}_{x+t}$) is fixed and deterministic, we find for $\bar{V}_{t}$,
$\bar{V}_{t}$ $=\mathbb{E}^{\mathbb{Q}}\left[\int_{t}^{T}e^{-\int_{t}^{s}r_{u}du}{}_{s-t}\bar{p}^{*}_{x+t}\,\widetilde{H}(A_{s},G^{I}_{s})\,ds\ \big|\ \mathcal{H}_{t}\right]$
$=\int_{t}^{T}{}_{s-t}\bar{p}^{*}_{x+t}\,\mathbb{E}^{\mathbb{Q}}\left[e^{-\int_{t}^{s}r_{u}du}\,\widetilde{H}(A_{s},G^{I}_{s})\ \big|\ \mathcal{H}_{t}\right]ds$
$=\int_{t}^{T}{}_{s-t}\bar{p}^{*}_{x+t}\,C_{t}(s,G^{I}_{s})\,ds\,$, (A.2)
where $C_{t}(s,G^{I}_{s})$ is given by (4.15) using again Assumption 4.2.
(ii) Lower bound. To minimize $V_{t}^{\mu}$ under (4.6) on each $(\lceil t\rceil+j,\lceil t\rceil+j+1]$, the same monotonicity argument shows that the conditional expected accumulated discounted payoff over $(\lceil t\rceil+j,\tau]$ is minimized by taking $\tau$ as small as possible. Since $\mu_{x,c}$ is continuous and takes value in a compact set, the infimum corresponds to $\underline{\mu\mkern-4.0mu}\mkern 4.0mu^{*}_{x,c}\equiv 0$ and $\tau=\lceil t\rceil+j$. However, $\lceil t\rceil+j\notin(\lceil t\rceil+j,\lceil t\rceil+j+1]$, hence no minimizer exists in the admissible set $\mathcal{T}_{(\lceil t\rceil+j,\lceil t\rceil+j+1]}$. Instead, for each $j=0,\ldots,\lfloor T\rfloor-\lceil t\rceil-1$, let $\tau_{j}^{(n)}\in\mathcal{T}_{(\lceil t\rceil+j,\lceil t\rceil+j+1]}$ be a sequence such that $\tau_{j}^{(n)}\downarrow\lceil t\rceil+j$ a.s. Define admissible controls $\mu^{(n)}\in\mathcal{W}$ by setting $\mu_{x,c}^{(n)}\equiv 0$ and placing the discrete jump on $(\lceil t\rceil+j,\lceil t\rceil+j+1]$ at time $\tau_{j}^{(n)}$ with size $\mu^{(n)}_{x,j}=1-\hat{p}_{j}$ for each $j$. Then the corresponding survival probabilities $_{s-t}p^{\mu^{(n)}}_{x+t}$ decrease pointwise for $n\to\infty$ to the deterministic limit
$_{s-t}\underline{p}^{*}_{x+t}=\prod_{j:\,\lceil t\rceil+j<s}\hskip-7.11317pt(1-\mu_{x,j})=\prod_{j:\,\lceil t\rceil+j<s}\hat{p}_{j}\,.$
Then, from equation (4.16), we write
$V^{\mu^{(n)}}_{t}=\mathbb{E}^{\mathbb{Q}}\left[\int_{t}^{T}e^{-\int_{t}^{s}r_{u}\,du}\hskip-5.69054pt\prod_{j\,:\,t<\tau^{(n)}_{j}\leq s}\hskip-5.69054pt(1-\mu_{x,j})\,\widetilde{H}(A_{s},G^{I}_{s})\,ds\ \big|\ \mathcal{H}_{t}\right].$
By dominated convergence (using $\mu_{x,j}\in[0,1]$ and $\mathbb{E}^{\mathbb{Q}}[\int_{0}^{T}|\widetilde{H}(A_{s},G_{s}^{I})|ds]<\infty$), we have as $n\to\infty$,
$V_{t}^{\mu^{(n)}}\downarrow\mathbb{E}^{\mathbb{Q}}\!\left[\int_{t}^{T}e^{-\int_{t}^{s}r_{u}\,du}\,{}_{s-t}\underline{p}^{*}_{x+t}\,\widetilde{H}(A_{s},G_{s}^{I})\,ds\,\Big|\,\mathcal{H}_{t}\right]=:\underline{V}_{t}.$
Finally, since $_{s-t}\underline{p}^{*}_{x+t}$ is deterministic, we obtain as in (i),
$\underline{V}_{t}=\int_{t}^{T}{}_{s-t}\underline{p}^{*}_{x+t}\,C_{t}(s,G_{s}^{I})\,ds,$
which concludes the proof.
A.3. Proof of Proposition 4.5
Proof.
(i) Upper bound. Fix the interval $(\lceil t\rceil+j,\lceil t\rceil+j+1]$ for some $j=0,\ldots,\lfloor T\rfloor-\lceil t\rceil-1$. Let again $\mathcal{T}_{(a,b]}$ denote the set of $\mathbb{H}$-stopping times on $(a,b]$. Therefore, for any $\tau\in\mathcal{T}_{(\lceil t\rceil+j,\lceil t\rceil+j+1]}$, we can rewrite the expected discounted payoff at time of death $\tau$ by
$\mathbb{E}^{\mathbb{Q}}\left[e^{-\int_{t}^{\tau}r_{s}ds}\,\widehat{H}(A_{\tau},G^{D}_{\tau})\ \big|\ \mathcal{H}_{t}\right]$ (A.3)
$=\mathbb{E}^{\mathbb{Q}}\left[e^{-\int_{t}^{\tau}r_{s}ds}\left(G^{D}_{\tau}+(A_{\tau}-G^{D}_{\tau})_{+}\right)\big|\ \mathcal{H}_{t}\right]$
$=\mathbb{E}^{\mathbb{Q}}\left[e^{-\int_{t}^{\tau}r_{s}ds}\,G^{D}_{\tau}\big|\ \mathcal{H}_{t}\right]+\mathbb{E}^{\mathbb{Q}}\left[e^{-\int_{t}^{\tau}r_{s}ds}(A_{\tau}-G^{D}_{\tau})_{+}\big|\ \mathcal{H}_{t}\right].$
Denoting $\widetilde{P}(t,s):=\mathbb{E}^{\mathbb{Q}}\left[e^{-\int_{t}^{s}r_{u}du}G^{D}_{s}\big|\ \mathcal{H}_{t}\right]=P(t,s)G^{D}_{s}=A_{0}P(t,s)e^{r_{g}s}$, we have under the assumption $r_{g}\geq f(t,s)$ for all $s\in[t,T]$, that $\widetilde{P}(t,s)$ is increasing in $s$. This is indeed direct from
$\partial_{s}\widetilde{P}(t,s)=P(t,s)A_{0}e^{r_{g}s}(-f(t,s)+r_{g})\geq 0\,$, (A.4)
since $\partial_{s}P(t,s)=-f(t,s)P(t,s)$. Moreover, under assumption (4.25), it is shown in Battauz and Rotondi (2022) that early exercise in an American call is never optimal and hence,
$\operatorname*{arg\,max}_{\tau\in\mathcal{T}_{(\lceil t\rceil+j,\lceil t\rceil+j+1]}}\mathbb{E}^{\mathbb{Q}}\left[e^{-\int_{t}^{\tau}r_{s}ds}(A_{\tau}-G^{D}_{\tau})_{+}\big|\ \mathcal{H}_{t}\right]=\lceil t\rceil+j+1\,.$ (A.5)
From the two statements above, we have shown that $s\mapsto C_{t}(s,G^{D}_{s}):=\mathbb{E}^{\mathbb{Q}}[e^{-\int_{t}^{s}r_{u}du}\,\widehat{H}(A_{s},G^{D}_{s})\mid\mathcal{H}_{t}]$ is nondecreasing on each year interval $(\lceil t\rceil+j,\lceil t\rceil+j+1]$. Therefore, using the same monotonicity argument (A.1) as in Proposition 4.22, we conclude that $\bar{\tau}^{*}_{j}=\lceil t\rceil+j+1$ and $\bar{\mu}^{*}_{x,c}(t)\equiv 0$. The life-table constraint (4.6) gives us directly the associated intervention size $\bar{\mu}^{*}_{x,j}=1-\hat{p}_{j}$. Since $\bar{\mu}^{*}$ is deterministic and from Assumption 4.2, we obtain the following upper bound
$\bar{V}_{t}$ $=\mathbb{E}^{\mathbb{Q}}\left[\sum_{j=0}^{\lfloor T\rfloor-\lceil t\rceil-1}e^{-\int_{t}^{\tau^{*}_{j}}r_{u}\,du}\left(\prod_{i<j}(1-\mu^{*}_{x,i})\right)\mu^{*}_{x,j}\,\widehat{H}(A_{\tau^{*}_{j}},G^{D}_{\tau^{*}_{j}})\,\Big|\,\mathcal{H}_{t}\right]$
$=\sum_{j=0}^{\lfloor T\rfloor-\lceil t\rceil-1}C_{t}(\lceil t\rceil+j+1,G^{D}_{\lceil t\rceil+j+1})\,\mu^{*}_{x,j}\,\prod_{i<j}\,(1-\mu^{*}_{x,i})\,$
$=\sum_{j=0}^{\lfloor T\rfloor-\lceil t\rceil-1}C_{t}(\lceil t\rceil+j+1_{j},G^{D}_{\lceil t\rceil+j+1})\,(1-\hat{p}_{j})\,\prod_{i<j}\,\hat{p}_{i}\,.\vskip 5.69054pt$
(ii) Lower bound.   The lower bound $\underline{V\mkern-4.0mu}\mkern 4.0mu_{t}$ follows directly from Proposition 4.22 (ii) and from the fact that $s\mapsto C_{t}(s,G^{D}_{s}):=\mathbb{E}^{\mathbb{Q}}[e^{-\int_{t}^{s}r_{u}du}\,\widehat{H}(A_{s},G^{D}_{s})\mid\mathcal{H}_{t}]$ is nondecreasing on each year interval $(\lceil t\rceil+j,\lceil t\rceil+j+1]$ as shown above in A.3 (i). ∎
A.4. Proof of Proposition 5.2
Proof.
(Step 1). First, we prove the equivalence between the original problem and the dual problem. For $\mu\in\mathcal{A}_{bounded}^{\delta}$, we have that, for any $\boldsymbol{\lambda}\in\mathbb{R}^{n}$,
$\mathbb{E}^{\mathbb{Q}}\Bigg[\,e^{-\int_{0}^{T}(r_{u}+\mu_{x}(u))du}H(A_{T},G_{T}^{A})+\int_{0}^{T}e^{-\int_{0}^{s}(r_{u}+\mu_{x}(u))du}\left(\widetilde{H}(A_{s},G_{s}^{I})+\mu_{x}(s)\widehat{H}(A_{s},G_{s}^{D})\right)ds\Bigg]$
$=$ $\mathbb{E}^{\mathbb{Q}}\Bigg[\,e^{-\int_{0}^{T}(r_{u}+\mu_{x}(u))du}H(A_{T},G_{T}^{A})+\int_{0}^{T}e^{-\int_{0}^{s}(r_{u}+\mu_{x}(u))du}\left(\widetilde{H}(A_{s},G_{s}^{I})+\mu_{x}(s)\widehat{H}(A_{s},G_{s}^{D})\right)ds\Bigg]$
$-\sum_{j=1}^{n}\lambda_{j}\left[\mathbb{E}^{\mathbb{P}}\left(e^{-\int_{0}^{j}\mu_{x}(s)ds}\right)-{}_{j}\hat{p}_{x}\right].$
Hence, we deduce that
$\sup_{\mu_{x}\in\mathcal{A}^{\delta}_{bounded}}\mathbb{E}^{\mathbb{Q}}\Bigg[\,e^{-\int_{0}^{T}(r_{u}+\mu_{x}(u))du}H(A_{T},G_{T}^{A})+\int_{0}^{T}e^{-\int_{0}^{s}(r_{u}+\mu_{x}(u))du}\left(\widetilde{H}(A_{s},G_{s}^{I})+\mu_{x}(s)\widehat{H}(A_{s},G_{s}^{D})\right)ds\Bigg]$
$\leq$ $\inf_{\boldsymbol{\lambda}\in\mathbb{R}^{n}}\hskip-1.42262pt\sup_{\mu\in\tilde{\mathcal{A}}^{\delta}_{bounded}}\hskip-5.12149pt\mathbb{E}^{\mathbb{Q}}\Bigg[e^{-\int_{0}^{T}(r_{u}+\mu_{x}(u))du}H(A_{T},G_{T}^{A})+\hskip-1.42262pt\int_{0}^{T}\hskip-3.55658pte^{-\int_{0}^{s}(r_{u}+\mu_{x}(u))du}\hskip-1.13809pt\left(\widetilde{H}(A_{s},G_{s}^{I})+\mu_{x}(s)\widehat{H}(A_{s},G_{s}^{D})\right)\hskip-1.13809ptds\Bigg]$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ -\sum_{j=1}^{n}\lambda_{j}\left[\mathbb{E}^{\mathbb{P}}\left(e^{-\int_{0}^{j}\mu_{x}(s)ds}\right)-{}_{j}\hat{p}_{x}\right].$
For $\boldsymbol{\lambda}\in\mathbb{R}^{n}$, let us define the inner functional $V(\boldsymbol{\lambda})$ as
$\hskip-2.84526ptV(\boldsymbol{\lambda})\hskip-1.42262pt=\hskip-5.406pt\sup_{\mu_{x}\in\tilde{\mathcal{A}}^{\delta}_{bounded}}$ $\mathbb{E}^{\mathbb{Q}}\Bigg[e^{-\int_{0}^{T}(r_{u}+\mu_{x}(u))du}H(A_{T},G_{T}^{A})\hskip-1.42262pt+\hskip-1.42262pt\int_{0}^{T}\hskip-4.2679pte^{-\int_{0}^{s}(r_{u}+\mu_{x}(u))du}\hskip-1.42262pt\left(\widetilde{H}(A_{s},G_{s}^{I})+\mu_{x}(s)\widehat{H}(A_{s},G_{s}^{D})\right)\hskip-1.42262ptds\Bigg]$ (A.6)
$-\sum_{j=1}^{n}\lambda_{j}\mathbb{E}^{\mathbb{P}}\left(e^{-\int_{0}^{j}\mu_{x}(s)ds}\right)$
As the set of admissible controls $\tilde{\mathcal{A}}^{\delta}_{bounded}$ is convex and closed, we deduce that there exists an optimal control $\mu^{*}(.;\boldsymbol{\lambda})\in\tilde{\mathcal{A}}^{\delta}_{bounded}$ such that
$V(\boldsymbol{\lambda}):=$ $\mathbb{E}^{\mathbb{Q}}\Bigg[\,e^{-\int_{0}^{T}(r_{u}+\mu_{x}^{*}(u;\boldsymbol{\lambda}))du}H(A_{T},G_{T}^{A})+\hskip-1.42262pt\int_{0}^{T}\hskip-4.2679pte^{-\int_{0}^{s}(r_{u}+\mu_{x}^{*}(u;\boldsymbol{\lambda}))du}\hskip-0.99585pt\left(\widetilde{H}(A_{s},G_{s}^{I})+\mu_{x}^{*}(s;\boldsymbol{\lambda})\widehat{H}(A_{s},G_{s}^{D})\right)\hskip-0.7113ptds\Bigg]$
$-\sum_{j=1}^{n}\lambda_{j}\mathbb{E}^{\mathbb{P}}\left(e^{-\int_{0}^{j}\mu_{x}^{*}(s;\boldsymbol{\lambda})ds}\right).$
Moreover, the functional $\boldsymbol{\lambda}\mapsto\left(V(\boldsymbol{\lambda})+\sum_{j=1}^{n}\lambda_{j}{}_{j}\hat{p}_{x}\right)$ is convex, there exists $\boldsymbol{\lambda}^{*}\in\mathbb{R}^{n}$ such that
$\inf_{\boldsymbol{\lambda}>0}V(\boldsymbol{\lambda})+\sum_{j=1}^{n}\lambda_{j}{}_{j}\hat{p}_{x}$ $=\min_{\boldsymbol{\lambda}>0}V(\boldsymbol{\lambda})+\sum_{j=1}^{n}\lambda_{j}{}_{j}\hat{p}_{x}=V(\boldsymbol{\lambda}^{*})+\sum_{j=1}^{n}\lambda_{j}^{*}{}_{j}\hat{p}_{x},$
and using the first order condition (FOC), we also have that $\boldsymbol{\lambda}^{*}$ is such that [^9] Since for all $\boldsymbol{\lambda}\in\mathbb{R}^{n}$, $\mu^{*}(\boldsymbol{\lambda})\in\tilde{\mathcal{A}}^{\delta}_{bounded}$, we know that, for all $t\leq T$, $(\mu_{x,a}(t)-\delta)_{+}\leq\mu_{x}^{*}(t;\boldsymbol{\lambda})\leq\mu_{x,a}(t)+\delta$, then, for j=1,,nj=1,...,n, $\inf_{\boldsymbol{\lambda}\in\mathbb{R}^{n}}\mathbb{E}^{\mathbb{P}}\left(e^{-\int_{0}^{j}\mu_{x}^{*}(t;\boldsymbol{\lambda})dt}\right)\leq{}_{j}\hat{p}_{x}\leq\sup_{\boldsymbol{\lambda}\in\mathbb{R}^{n}}\mathbb{E}^{\mathbb{P}}\left(e^{-\int_{0}^{j}\mu_{x}^{*}(t;\boldsymbol{\lambda})dt}\right)$ and by continuity of the functional, we deduce the the existence of $\boldsymbol{\lambda}^{*}\in\mathbb{R}^{n}$ that satisfies the FOC.
$\mathbb{E}^{\mathbb{P}}\left(e^{-\int_{0}^{j}\mu_{x}^{*}(s;\boldsymbol{\lambda}^{*})ds}\right)={}_{j}\hat{p}_{x},\quad j=1,...,n.$ (A.7)
Since $\mu_{x}^{*}(.;\boldsymbol{\lambda}^{*})$ satisfies (A.7), we have that $\mu_{x}^{*}(.;\boldsymbol{\lambda}^{*})\in\mathcal{A}_{bounded}^{\delta}$ and thus the solution of the original problem is such that
$\hskip-5.69054pt\sup_{\mu_{x}\in\mathcal{A}^{\delta}_{bounded}}\hskip-4.2679pt\mathbb{E}^{\mathbb{Q}}\Bigg[\,e^{-\int_{0}^{T}(r_{u}+\mu_{x}(u)du}H(A_{T},G_{T}^{A})\hskip-0.7113pt+\hskip-1.42262pt\int_{0}^{T}\hskip-4.2679pte^{-\int_{0}^{s}(r_{u}+\mu_{x}(u))du}\hskip-0.99585pt\left(\widetilde{H}(A_{s},G_{s}^{I})+\mu_{x}(s)\widehat{H}(A_{s},G_{s}^{D})\right)\hskip-0.99585ptds\Bigg]$
$\hskip 4.2679pt=\min_{\boldsymbol{\lambda}>0}V(\boldsymbol{\lambda})+\sum_{j=1}^{n}\lambda_{j}{}_{j}\hat{p}_{x}$
$\hskip 4.2679pt=\mathbb{E}^{\mathbb{Q}}\Bigg[e^{-\int_{0}^{T}(r_{u}+\mu_{x}^{*}(u;\boldsymbol{\lambda}^{*}))du}H(A_{T},G_{T}^{A})\hskip-0.99585pt+\hskip-1.42262pt\int_{0}^{T}\hskip-4.2679pte^{-\int_{0}^{s}(r_{u}+\mu_{x}^{*}(u;\boldsymbol{\lambda}^{*}))du}\hskip-0.99585pt\left(\widetilde{H}(A_{s},G_{s}^{I})+\mu_{x}^{*}(s;\boldsymbol{\lambda}^{*})\widehat{H}(A_{s},G_{s}^{D})\right)\hskip-0.99585ptds\Bigg]$
(Step 2). To conclude to proof, it remains to prove that, for any $\boldsymbol{\lambda}\in\mathbb{R}^{n}$, $\bar{v}(0,A_{0},r_{0},Z_{0};\boldsymbol{\lambda})$ is equal to $V(\boldsymbol{\lambda})$ defined by (A.6). Since we assume that there exists $\bar{v}(t,a,r,z;\boldsymbol{\lambda})\in C^{1,2,2,2}$ solution to (5.18) that satisfies boundary conditions given by (5.16), then using classical results from dynamic programming (see for instance Touzi (2012)) or by adapting the arguments used in the proof of (Li and Szimayer, 2011, Theorem 1) in the context of an uncertainty mortality control problem, it follows that $\bar{v}(t,a,r,z;\boldsymbol{\lambda})$ solving the HJB equation solves the stochastic control problem (5) such that
$V(\boldsymbol{\lambda})=$ $\bar{v}(0,A_{0},r_{0},Z_{0};\boldsymbol{\lambda}),$
and the proof is complete. ∎
A.5. Proof of Proposition 5.4
Proof.
To prove the statement, we must show that
$\tilde{\mathcal{A}}=\bigcup_{\delta>0}\mathcal{A}^{\delta}_{bounded}.$
Obviously, we have that
$\bigcup_{\delta>0}\mathcal{A}^{\delta}_{bounded}\subseteq\tilde{\mathcal{A}}$.
Let us prove that
$\tilde{\mathcal{A}}\subseteq\bigcup_{\delta>0}\mathcal{A}^{\delta}_{bounded}.$
Formally, we have to prove that for all $\mu_{x}\in\tilde{\mathcal{A}}$, there exists $\delta>0$ such that $\mu_{x}\in\mathcal{A}^{\delta}_{bounded}$. For a given $\mu_{x}\in\tilde{\mathcal{A}}$, define
$Y(\omega):=\sup_{t\in[0,T]}\big|\mu_{x}(t;\omega)-\mu_{x,a}(t)\big|.$
Since we assume that $\mu_{x}\in\tilde{\mathcal{A}}$, then
$Y(\omega)<+\infty.$
Consequently, for $\mu_{x}\in\tilde{\mathcal{A}}$, we can choose
$\delta:=\operatorname{ess\,sup}_{\omega}Y(\omega)<\infty,$
which implies $\mu_{x}\in\mathcal{A}^{\delta}_{\mathrm{bounded}}$. Since $\mu_{x}$ was arbitrary, we conclude that
$\bigcup_{\delta>0}\mathcal{A}^{\delta}_{\mathrm{bounded}}=\tilde{\mathcal{A}}$.
Disclosure statement The authors declare that they have no conflict of interest.
Funding details This work was supported by the Swiss National Science Foundation (SNSF) under Grant no. 10003723. Edouard Motte is a FRIA grantee of the Fonds de la Recherche Scientifique - FNRS, Belgium.